### v1

# Who Does the Giant Number Pile Like Best: Analyzing Fairness in Hiring Contexts

Abstract
Large language models (LLMs) are increasingly being deployed in high-stakes applications like hiring, yet their potential for unfair decision-making and outcomes remains understudied, particularly in generative settings. In this work, we examine the fairness of LLM-based hiring systems through two real-world tasks: resume summarization and retrieval. By constructing a synthetic resume dataset and curating job postings, we investigate whether model behavior differs across demographic groups and is sensitive to demographic perturbations. Our findings reveal that race-based differences appear in approximately 10% of generated summaries, while gender-based differences occur in only 1%. In the retrieval setting, all evaluated models display non-uniform selection patterns across demographic groups and exhibit high sensitivity to both gender and race-based perturbations. Surprisingly, retrieval models demonstrate comparable sensitivity to non-demographic changes, suggesting that fairness issues may stem, in part, from general brittleness issues. Overall, our results indicate that LLM-based hiring systems, especially at the retrieval stage, can exhibit notable biases that lead to discriminatory outcomes in real-world contexts.
Who Does the Giant Number Pile Like Best: Analyzing Fairness in Hiring Contexts
Preethi Seshadri UC Irvine preethis@uci.edu                        Seraphina Goldfarb-Tarrant Cohere seraphina@cohere.com
1 Introduction
Large language models (LLMs) are increasingly being used for real-world, high-stakes domains, including medicine, law, and hiring. As LLMs are incorporated into critical decision-making processes, ensuring reliable and responsible usage is paramount. A key aspect of this effort includes anticipating and preventing specific risks and harms, such as identifying biases that could unfairly disadvantage certain demographic groups in hiring screening processes.
Prior work focuses on two types of harms: representational harms (i.e., representing certain groups in a less favorable light than others, demeaning them, or failing to recognize their existence altogether) and allocational harms (i.e., allocating resources or opportunities or unfairly to different social groups) (Barocas et al., 2017; Blodgett et al., 2020). There is a substantial body of work that analyzes representational harms in LLMs (Zhao et al., 2018; Abid et al., 2021; Kirk et al., 2021; Cheng et al., 2023; Gadiraju et al., 2023), yet allocational harms—which are the primary harm at play in high-stakes situations—remain understudied. Given the growing use of LLMs in consequential applications, it is especially important to examine whether models produce unfair outcomes or performance gaps between different groups (allocational harms, also referred to as allocational fairness).
Although there are a few studies on LLM harms in real-world settings (e.g., hiring) (Tamkin et al., 2023; An et al., 2024; Haim et al., 2024; Nghiem et al., 2024), there is minimal work closely simulating actual deployment scenarios. By using highly simplified setups and tasks, these prior works may fail to predict real-world outcomes and as a result provide an incomplete assessment of harms. Investigations into LLM harms should ensure ecological validity (Goldfarb-Tarrant et al., 2021; Cao et al., 2022); they should either be grounded in realistic scenarios and tasks that match how these systems are used in practice, or use a proxy that has been proven to be predictive of real world outcomes.
Additionally, these existing studies have cast their investigations of allocational harms in LLMs as discrete classification tasks (such as binary accept/reject decisions) or quantitative predictions (such as determining salary levels). This makes the implementation and evaluation more straightforward, but does not match the real use of generative LLMs (Kelly, 2023; Yin et al., 2024). In contrast, there is limited work on allocational harms in generative settings (Wan et al., 2023), like summarization, since measuring how generated text might create disparities is more open-ended and complex than analyzing classification predictions.
In this work, we examine whether LLMs are fair in real-world contexts for hiring. We choose hiring due to the growing adoption of LLMs in hiring tools (Sánchez-Monedero et al., 2020). Hiring is a high-stakes domain with clear societal risks – unfair systems may discriminate against already marginalized groups, risking further underrepresentation throughout the hiring pipeline (Dastin, 2018; Raghavan et al., 2020; Suresh and Guttag, 2021). We focus on two tasks that mirror how LLMs are integrated into hiring and recruitment workflows: resume summarization and resume retrieval with respect to a job post (Herman, 2024; Humanly, 2024). We then examine whether model behavior in these tasks is sensitive to gender and race demographic perturbations in resumes (e.g., name perturbations or tailored extracurricular content). Specifically, we ask the following questions:
  • RQ1: Do model generated summaries differ meaningfully across demographic groups?
  • RQ2: In a retrieval setting, do models disparately select resumes across demographic groups?
  • RQ3: In a retrieval setting, how sensitive are model rankings to demographic and non-demographic perturbations in resumes?
In this work, we: (1) create a synthetic resume dataset and curate a job post dataset, (2) propose metrics to measure fairness in generative and retrieval settings for two real-world hiring tasks, validated by an expert human preference study, and (3) perform a study investigating the fairness of an automated LLM hiring system based on real world usage, for 7 models (3 generative, 4 retrieval). [^1] The data and code can be found at https://github.com/preethisesh/hiring_fairness.
In the summarization setting, on average, we observe meaningful differences between summaries generated for different racial groups nearly 10% of the time, as opposed to only 1% for gender (RQ1). On the other hand, for retrieval, we find that all models are prone to non-uniformly selecting resumes across demographic groups (RQ2). In addition, all models produce rankings that are highly sensitive to both name and race demographic perturbations (RQ3). We also find that model sensitivity to non-demographic perturbations is notable and sometimes even on par with demographic perturbations (RQ3), suggesting that fairness issues may largely be due to model brittleness as a whole, rather than encoded biases only. Overall, our results demonstrate that an LLM hiring system with automated resume retrieval and summarization exhibits considerable bias, primarily stemming from the initial retrieval stage.
2 Methodology
To study fairness in hiring, we consider an LLM-based pipeline with two components: resume retrieval with respect to a job post (using an embedding model), and resume summarization (using an LLM). [^2] We analyze both components because they reflect real-world usage of automation to streamline hiring processes. We focus on summarization first because it is more neglected in research, though in a pipeline it would come after retrieval (as summarization would be of retrieved resumes). We investigate the impact of perturbing the demographic group of a resume on each component of the pipeline and look for: (1) systematic differences in generated resume summaries and (2) changes in similarity of a resume to a job posting, and as a result its ranking in a resume set.
Let $D$ represent a set of resumes, where each resume $d\in D$ has demographic label $l(d)$. For each resume, let $d^{\prime}$ represent any perturbed version of $d$ where $l(d^{\prime})\neq l(d)$. For each $d$, there can be several perturbed versions $d^{\prime}$ along various demographic dimensions (e.g., gender or race, to name a few). The resume content remains largely unchanged except for the demographic perturbation.
2.1 Summarization
Perturbed resumes are by design highly similar to original resumes, so we expect generated summaries for original and perturbed resumes to also maintain a high degree of similarity. In other words, we are testing for invariance, since we expect the output to change minimally after perturbing the input (Ribeiro et al., 2020).
In order to assess invariance, we need a way to measure whether the original and perturbed summaries differ meaningfully. For cost and scalability reasons, we rely on an automated approach—human preferences are expensive and cannot be collected frequently, especially if these preferences are sourced from HR staff. Additionally, collecting human preferences for a specific dataset does not extend to other analyses with new datasets.
Automated measures can evaluate whether specific properties of generated summaries differ, where differences in these properties would be undesirable in practice. For instance, summaries for a specific demographic group should not be written more positively than summaries for other demographic groups; this could lead to disparity in hiring outcomes. We use the following measures as proxies for undesirable variation that could influence the decision of an HR staff reading the summary: reading ease, reading time, polarity, subjectivity, and regard (Appendix A.5). We verify that these measures capture meaningful differences in practice by conducting a preference task annotated by HR staff. This study shows that these are good proxies for human preferences (Appendix A.6).
Fairness Metric
We compute an invariance metric that aggregates the proposed measures for both original and perturbed resume summaries. For each measure, we do a paired t-test between scores for original and perturbed summaries, repeated for each demographic perturbation. We then calculate how often the null hypothesis (that the mean difference between paired summaries is zero) is rejected, i.e., how often invariance is violated. We choose a significance level of $\alpha=0.05$, and apply Bonferroni correction to account for Type 1 errors with multiple comparisons (Bland and Altman, 1995).
[…]
2.2 Retrieval
In contrast to summarization, which relies on only the resume, retrieval utilizes both a query (i.e., job posting) and candidates to select (i.e., resumes). The top-$n$ of the resulting resumes sorted by similarity then make it to the next stage of the pipeline. Similar to the previous setting, we assume that demographically perturbing a resume should have no impact on its relevance to a job posting.
Given an embedding model $\mathcal{M}$ and a similarity measure $sim$, we compute the similarity between embeddings for resume $d$ and a job posting $p$. In practice, we use cosine similarity (also used in Wilson and Caliskan (2024)). We define the set $S(p)$, which represents the set of similarity values between each resume $d\in D$ and job posting $p$:
$S(p)=\{sim(\mathcal{M}(d),\mathcal{M}(p))\mid d\in D\}$
We transform the similarity values between each resume and job posting, $s_{i}$, into a rank such that lower ranks indicate higher similarity (note: the highest similarity value receives rank 1):
$\text{rank}_{p}(s_{i})=|\{s_{j}\in S(p)\mid s_{j}>s_{i}\}|+1$
Let $D_{n}(p)$ represent the set of top-$n$ resumes from $D$, which are the resumes with the $n$ lowest ranks (i.e., highest similarities) for job posting $p$:
$D_{n}(p)=\{d_{i}\in D\mid\text{rank}(s_{i})\leq n\}$
Fairness Metrics
We compute two fairness metrics for retrieval, non-uniformity, proposed by Wilson and Caliskan (2024), and exclusion, which we introduce below. Further intuition about these metrics are provided in Appendix A.7.
Non-uniformity
evaluates whether the top-$x$(where $x$ is a percentage instead of a fixed number $n$) of resumes are uniformly distributed with respect to the demographic groups being considered. To compute non-uniformity, we consider the set of top-$x$of resumes retrieved from the combined pool of all demographically perturbed versions, which we denote as $D_{x}^{\prime}(p)$. Following the approach from Wilson and Caliskan (2024), we perform a chi-squared goodness-of-fit test to determine whether the demographic makeup of $D_{x}^{\prime}(p)$ deviates from the uniform distribution.
Exclusion
evaluates how often resumes in the set of top-$n$ resumes are excluded (i.e., the ranking falls outside top-$n$) after perturbation. Ideally, $\mathcal{M}$ should be robust to demographic perturbations, yielding nearly identical similarity scores and rankings for both resume $d$ and demographically perturbed version $d^{\prime}$. Exclusion directly tests this behavior by measuring the extent to which the set of top-$n$ resumes differs after perturbation.
$\text{exclusion}_{n}(p)=\frac{|{\text{rank}_{p}(d^{\prime})>n\mid d\in D_{n}(p)}|}{|D_{n}(p)|}$
3 Experimental Setup
In this section, we provide an overview of the datasets, perturbations, and models used in our evaluation.
3.1 Data
Resumes
Resumes are sourced through social media platforms including LinkedIn, Slack, and X. Given the authors’ professional circles, the sample skews heavily toward tech and academic professionals. For privacy reasons, we anonymize resumes using Presidio anonymize, [^3] https://microsoft.github.io/presidio/anonymizer/ which masks PII entities. To further mitigate privacy concerns and enable us to release the dataset, we use the collected resumes as examples to generate synthetic resumes, instead of using original resumes in our analysis. We generate 525 resumes across 21 professions using Cohere’s Command-R model (Cohere, 2024). All synthetic resumes are free of explicit demographic information, until we add it as part of experimentation.
In addition to our curated resume set, we use a publicly available resume dataset from Kaggle (Bhawal, 2021). [^4] Has a CC0: Public Domain License These resumes differ in two notable ways: (1) the Kaggle resumes are less structured and formatted than synthetically generated resumes, and (2) they include a more diverse set of fields (e.g., apparel, construction, fitness, etc.). We sample 1175 Kaggle resumes across 24 fields.
Job Posts
Our resume dataset consists of two types: realistic synthetic resumes generated for specific roles (e.g., data analyst) and actual resumes from Kaggle labeled with broader field categories (e.g., construction). For each profession/field (we choose 11 from generated resumes and 11 from Kaggle resumes), we carefully select 7 detailed LinkedIn job postings that clearly describe the role and outline qualifications. For Kaggle resumes, we match postings to roles within each broader field (e.g., choosing “construction manager” postings for the construction field). In total, we collect 154 job postings across different professions.
3.2 Demographic Perturbations
We use names as a proxy for gender and racial information. Initially, all resumes are free of names. We then add names using the curated set of names from Yin et al. (2024). [^5] Uses voter registration data from North Carolina to identify demographically-distinct names. We consider four demographic groups, each with 100 unique names: Black female (FB), White female (FW), Black male (MB), and White male (MW). Following Wilson and Caliskan (2024), we use the demographically neutral “Williams” as the last name for all groups and only vary first name.
In actual resumes, demographic information can be encoded in more than just the name. Therefore, for synthetically generated resumes, we perform an additional augmentation step that adds extracurricular information to the resumes (similar to Glazko et al. (2024)). We use Command-R to augment resumes by including awards, clubs and leadership, and mentorship and volunteering experiences that are reflective of the individual’s background and identity (more details in Appendix A.3 and A.4). In addition to explicit gender and race information from names, extracurricular information can reinforce demographic signal by providing further explicit and implicit cues.
(a) GPT-4o
(b) Command-R
(c) Mixtral 8x7b
Figure 1: Summarization Results: Percentage of t-tests for which we reject null hypothesis (i.e., invariance is violated), separated by completion model and perturbation.
3.3 Non-Demographic Perturbations
We do two non-demographic perturbation experiments for retrieval to assess baseline sensitivity of embedding models that is not due to demographics.
Within Group Name Perturbations
As a baseline comparison to performing name perturbations between different demographic groups (e.g., White female $\rightarrow$ Black Female), we assess whether models are sensitive to same group demographic perturbations (e.g., White female $\rightarrow$ White Female). By doing so, we disentangle how much bias is due demographic changes vs. overall model sensitivity to name changes.
To control for the effects of frequency (Ethayarajh et al., 2019), we bin names in each demographic group according to their frequency in the Pile dataset (Gao et al., 2020), and match names based on the bin. Therefore, name perturbations should result in a similarly occurring name within the same demographic group.
Non-Name Perturbations
We also assess whether model rankings are sensitive to non-name perturbations. This allows us to examine whether models lack robustness more broadly, or mainly exhibit brittleness with respect to name changes.
We test two perturbation types: (1) random character swapping, which does not impact readability or comprehension in a meaningful way and (2) replacing new lines in the resume with a single space instead, which targets formatting without modifying content. For character swapping, we choose 10 random characters in the resume and swap based on neighboring keys to simulate typos.
3.4 Models
Summarization
We generate summaries using a combination of closed and open state-of-the-art language models: GPT-4o (OpenAI, 2024), Command-R (Cohere, 2024), and Mixtral 8x7B (Jiang et al., 2024). For summary instructions, we vary the generation length (100 and 200 words) as well as the point of view (first and third person) specified in the prompt. We generate summaries with temperatures of 0.0 and 0.3.
Retrieval
For retrieval, we select four popular dense embedding models used in retrieval augmented generation (RAG) systems (Lewis et al., 2020): OpenAI’s text-embedding-3-small and text-embedding-3-large models, Cohere’s embed-english-v3.0, and Mistral’s mistral-embed.
(a) Top-5%
(b) Top-10%
Figure 2: Non-uniformity metric for top-5 and top-10% of retrieved resumes. Separated (sep) measures the % of individual job posts where the top-$x$of resumes are non-uniformly distributed, while pooled (pool) measures the % of occupations where the top-$x$of resumes across job posts for that occupation are non-uniformly distributed.
4 Results
In this section, we evaluate the use of LLMs in two real-world hiring tasks: resume summarization and retrieval. Unless otherwise mentioned, we present results for generated resumes below, and include results for Kaggle resumes in the Appendix. Similar trends and findings hold for both datasets.
4.1 Summarization
We analyze whether generated summaries differ meaningfully across demographic groups (RQ1) by examining invariance violations, i.e., the percentage of t-tests that yield significant differences in our automated measures. We measure across summaries with different characteristics (length, point of view, and temperature). Figure 1 displays results for generated resumes, grouped by completion model and demographic perturbation.
Mixtral 8x7B exhibits the highest average invariance violations (8.59%), followed by GPT-4o (3.91%) and Command-R (2.74%). All models violate invariance much more for resumes that differ by race vs. gender. For example, the average invariance violations for race perturbations is 15.62% for Mixtral 8x7B, as compared to only 1.56% for gender. In summary, we observe that models exhibit some but not considerable discrepancies between generated summaries for different demographic groups, with different models having potentially large differences in sensitivity.
4.2 Retrieval
We now shift to a retrieval setting and ask: do models exhibit fairness issues in resume retrieval? Our analysis tackles this question from both a distributional (RQ2) and robustness (RQ3) perspective.
4.2.1 Non-Uniformity
Do models disparately select resumes across demographic groups?
We compute non-uniformity (i.e., how often top retrieved resumes have non-uniform demographic distributions). Non-uniformity differs considerably across models, choice of top-$x$ percent, and whether or not resumes are pooled across occupations (Figure 2).
(a) $n=5$
(b) $n=10$
(c) $n=100$
Figure 3: Exclusion metric for retrieval after performing gender and race name perturbations for the top-5, top-10, and top-100 retrieved resumes. Lower values indicate models are less sensitive to demographic perturbations.
We observe that embed-english-v3.0 exhibits the highest non-uniformity on average, where 6.90% of job posts and 45.45% of occupations have non-uniformly distributed resumes. We see that increasing top-$x$ from 5% to 10% and pooling resumes both yield higher non-uniformity across all models. In particular, pooling resumes across occupations can produce massive changes – for example, we see that on average across models, non-uniformity goes from 3.66% $\rightarrow$ 30.68% after pooling, resulting in a nearly 750% increase. However, this reflects a sensitivity in the metric itself more than a change in the shape of the distribution. Increasing top-$x$ and pooling both increase sample size, which can lead to rejecting the null hypothesis that the distribution is uniform in cases where the null hypothesis previously failed to be rejected.
Different models demonstrate distinct patterns of bias. For example, when considering the highest represented groups in the top-10% of resumes with respect to each job post, White females are the top group in 37 job posts (48.05%), while Black males are the top group in only 3 job posts (3.90%) for embed-english-v3.0. In contrast, for mistral-embed, White males are the top group in 56 job posts (72.73%), while White females are the top group in only 4 job posts (5.19%). We cannot determine why such differences exist without access to datasets or training details, but it is notable and somewhat surprising that models do not uniformly prefer one demographic group over the other.
4.2.2 Exclusion
How sensitive are models to gender and race perturbations?
We compute exclusion (i.e., how often top retrieved resumes are excluded from the set of top-$n$ resumes after demographic perturbations). We find that all models display notable sensitivity to gender and race name perturbations, as shown in Figure 3. When considering top-5 resumes, we find that models tend to exclude perturbed resumes nearly half the time (45.75% on average).
Across both gender and race name perturbations, and different $n$ values, text-embedding-3-small and text-embedding-3-large have the highest exclusion, while embed-english-v3.0 consistently has the lowest exclusion. As expected, exclusion lowers as $n$ increases, since larger $n$ values are less restrictive and consider a larger set of retrieved resumes. That being said, exclusion for $n=100$ is still considerable, as all models have exclusion > 12%. [^6] In practice we expect $n$ to be low for filtering candidates.
For exclusion, models have similar sensitivity to both gender and race perturbations. Overall, average exclusion for gender is 31.78% on generated resumes (25.47% on Kaggle resumes) vs. 31.66% on generated resumes (25.40% on Kaggle resumes) for race. [^7] Kaggle resumes exhibit similar patterns but lower exclusion magnitude than generated resumes. This is likely because generated resumes are tech-focused and closer in content. These results contrast our summarization findings, where models show greater invariance violations on race vs. gender perturbations. Our analysis reveals that the set of top retrieved resumes with respect to a job posting is highly brittle, as merely altering the demographic on resumes (through names only) often results in otherwise identical resumes dropping out of the top-$n$ results.
(a) Gender, $n=5$
(b) Race, $n=5$
Figure 4: Directional differences in exclusion metric for retrieval after applying name perturbations (i.e., separating based on perturbation direction). M→F perturbs male to female names and F→M perturbs female to male names, while W→B perturbs White to Black names and B→W perturbs Black to White names.
Does model sensitivity to perturbations differ based on the direction of perturbation?
We can also partition the results based on direction of perturbation (Figure 4). We find that the magnitude of difference exceeds a threshold of 5% (e.g., the absolute difference in exclusion for M $\rightarrow$ F vs. F $\rightarrow$ M or W $\rightarrow$ B vs. B $\rightarrow$ W for the same model, $n$, and resume type is $>5$) in 26/48 cases across generated and Kaggle resumes, indicating models often exhibit higher sensitivity to one direction of perturbation over the other. In particular, the gender directional difference is notable for mistral-embed, going from 63.28% for M $\rightarrow$ F to 27.93% for F $\rightarrow$ M, for generated resumes with $n=5$ (Figure 4).
(a) Gender, $n=5$
(b) Race, $n=5$
Figure 5: Deltas (differences) in exclusion metric for retrieval after performing demographic perturbations with names as well as extracurricular information as compared to only names, separated by direction. As expected, adding extracurricular information increases sensitivity to perturbations.
We also observe that models exhibit opposite directional trends for gender and race in Figure 4. For gender, all models except mistral-embed are more sensitive when perturbing female names (marginalized) to male names (non-marginalized); for race, all models except mistral-embed are more sensitive when perturbing White names (non-marginalized) to Black names (marginalized). These results highlight an asymmetry in how models handle different types of demographic changes.
Are models more sensitive when perturbing both names and extracurricular information, as opposed to names only?
Figure 5 shows that, as expected, models tend to be more sensitive when perturbing extracurricular information (e.g., awards, clubs, volunteering, etc.) in addition to names. On average, across models and $n$ values, we observe the following increases in exclusion: 9.35% for M $\rightarrow$ F, 8.06% for F $\rightarrow$ M, 16.41% for W $\rightarrow$ B, and 2.90% for B $\rightarrow$ W.
For gender, adding extracurricular information results in comparable increases in exclusion for both directions. In contrast, adding extracurricular information for race results in highly asymmetric increases in exclusion. W $\rightarrow$ B averages more than 5x the increase of B $\rightarrow$ W changes. We observe that adding extracurricular information results in non-uniform increases to exclusion, which suggests that models may encode and utilize various types of demographic signal differently. This finding is notable given that prior work often examines a single way of encoding demographics, overlooking how various signals interact and compound.
More broadly, do models exhibit brittleness to non-demographic perturbations?
We consider two sets of perturbation analyses that are non-demographic: (1) How sensitive are models to within-group name perturbations? and (2) How sensitive are models to non-name perturbations? Even when perturbing names within the same demographic group, models surprisingly exhibit similar levels of sensitivity to those observed with demographic name perturbations (Figure 6(a)).
(a) Within Group Name Perturbation
(b) Non-name Perturbation
Figure 6: Exclusion metric for retrieval after performing non-demographic perturbations (i.e., within group name changes - left, and modifying spacing and adding typos - right). Results shown for $n=5$.
For non-name perturbations, we find that models are extremely sensitive to both spacing and typos, although to a lesser extent than name perturbations. As shown in Figure 6(b), most models demonstrate higher sensitivity to spacing than typos, though there is surprising sensitivity to both. In particular, mistral-embed excludes resumes from the top-5 set 72.76% of the time solely based on spacing, which indicates that simple formatting choices can have a massive impact on fairness (in this case, much more than names). Generated resumes display nearly twice the sensitivity to spacing changes compared to web-scraped Kaggle resumes (27.98% vs. 15.35% averaged across models and $n$ values), likely due to their more structured and formatted nature. In summary, we observe that retrieval models lack robustness more generally, which might contribute to fairness-related issues.
5 Discussion
Our results highlight important issues about model fairness and brittleness in real-world contexts. First, we observe that model rankings in a retrieval setting are impacted considerably by both subtle demographic and non-demographic changes. In practice, these differences would lead to unintended exclusion, with candidates being eliminated from consideration at an early stage. We also observe that subtle demographic differences in resumes can alter the way candidates are discussed in generated summaries. As a result, even candidates who make it past the initial resume screening stage may be portrayed differently based on demographic attributes. These differences can impact downstream decision-making, since recruiters might rely solely on summaries to shortlist candidates.
Additionally, it is important to consider compositional effects when combining components (i.e., ranking $\rightarrow$ summarization), since biases can compound due to the sequential nature of these tasks. Therefore, the candidate pool may (1) leave out qualified candidates through retrieval bias, as well as (2) differentially represent candidates through summarization bias. Applying this to our results, at worst case, we find that modifying racial indicators on resumes from W $\rightarrow$ B (using both names and extracurricular information) can result in roughly (1) 70% of candidates being filtered out at the resume screening stage, and (2) 15% of remaining candidates being depicted in less preferable ways.
Our analysis also reveals that all models are sensitive to non-demographic perturbations. These findings suggest that model unfairness may partially stem from more general robustness issues, rather than solely from encoded biases. While these insights do not change the real-world impacts of discriminatory outcomes, understanding that disparate treatment can arise from multiple factors—including but not limited to undesirable biases—provides a more complete and nuanced picture for addressing fairness issues. Moreover, given that retrieval models are commonly used in RAG systems, these issues may extend to various settings and applications beyond HR.
6 Conclusion
We examine fairness in LLM-based hiring systems by analyzing two key components: applicant ranking and summary generation. We propose two new metrics to measure fairness in retrieval and generative settings, and find that a hiring pipeline with these components produces biased outcomes, particularly during the initial retrieval phase. In addition, we see unexpected sensitivity to minor non-demographic changes, indicating a lack of overall robustness that may contribute to unfair outcomes. We hope that our work underscores the need for more real-world evaluations of LLM harms, especially in other high-stakes domains.
Limitations
Our analysis focuses exclusively on English resumes and job posts, which poses limitations. Future research should investigate fairness considerations in multilingual settings and examine whether our conclusions hold across various languages. Additionally, cultural norms likely influence how candidates present themselves and describe their professional experience, qualifications, and achievements. Understanding these nuances is crucial for evaluating and developing hiring systems that serve diverse global talent pools.
While our analysis examines whether hiring systems behave differently for various gender (male and female) and racial (White and Black) groups, it is meant to be illustrative rather than exhaustive and only covers a subset of gender and racial identities. We only consider binary gender biases, and exclude non-binary gender biases from our analysis (e.g., if a candidate explicitly declares pronouns on their resume). We also only focus on Black and White racial groups, since this is a common emphasis in fairness studies. In addition, future research should expand beyond these commonly investigated biases and analyze to what extent other types of demographic information (e.g., age and nationality) impact LLM fairness in hiring.
Appendix A Appendix
A.1 Names
We use White male, Black male, White female, and Black female names curated by Yin et al. (2024), which we list below:
White male
Adam, Aidan, Aiden, Alec, Andrew, Austin, Bailey, Benjamin, Blake, Braden, Bradley, Brady, Brayden, Brendan, Brennan, Brent, Bret, Brett, Brooks, Carson, Carter, Chad, Chase, Clay, Clint, Cody, Colby, Cole, Colin, Collin, Colton, Conner, Connor, Conor, Cooper, Dalton, Davis, Dawson, Dillon, Drew, Dustin, Dylan, Eli, Ethan, Gage, Garrett, Graham, Grant, Grayson, Griffin, Harley, Hayden, Heath, Holden, Hunter, Jack, Jackson, Jacob, Jake, Jakob, Jeffrey, Jody, Jon, Jonathon, Kurt, Kyle, Landon, Lane, Liam, Logan, Lucas, Luke, Mason, Matthew, Max, Owen, Parker, Peyton, Philip, Randall, Reid, Riley, Ross, Scott, Seth, Shane, Skyler, Stuart, Tanner, Taylor, Todd, Tucker, Walker, Weston, Wyatt, Zachary, Zachery, Zackary, Zackery, Zane
Black male
Akeem, Alphonso, Amari, Antione, Antoine, Antwain, Antwan, Antwon, Cedric, Cedrick, Cornell, Cortez, Daquan, Darius, Darnell, Darrius, Dashawn, Davion, Davon, Davonte, Deandre, Deangelo, Dedrick, Demarcus, Demario, Demetrius, Demond, Denzel, Deonte, Dequan, Deshaun, Deshawn, Devante, Devonte, Dominique, Donnell, Donta, Dontae, Donte, Hakeem, Ishmael, Jabari, Jaheim, Jaleel, Jamaal, Jamal, Jamar, Jamari, Jamel, Jaquan, Javon, Jaylen, Jermaine, Jevon, Juwan, Kareem, Keon, Keshawn, Kevon, Keyon, Kwame, Lamont, Malik, Marques, Marquez, Marquis, Marquise, Mekhi, Montrell, Octavius, Omari, Prince, Raekwon, Raheem, Raquan, Rashaad, Rashad, Rashaun, Rashawn, Rasheed, Rico, Roosevelt, Savion, Shamar, Shaquan, Shaquille, Stephon, Sylvester, Tevin, Travon, Tremaine, Tremayne, Trevon, Tyquan, Tyree, Tyrek, Tyrell, Tyrese, Tyrone, Tyshawn
White female
Abby, Abigail, Aimee, Alexandra, Alison, Allison, Allyson, Amanda, Amy, Ann, Anna, Anne, Ashlyn, Bailey, Beth, Bethany, Bonnie, Brooke, Caitlin, Caitlyn, Cara, Carly, Caroline, Casey, Cassidy, Cassie, Claire, Colleen, Elisabeth, Elizabeth, Ellen, Emily, Emma, Erin, Ginger, Hailey, Haley, Hannah, Hayley, Heather, Heidi, Holly, Jaclyn, Jaime, Jeanne, Jenna, Jennifer, Jill, Jodi, Julie, Kaitlin, Kaitlyn, Kara, Kari, Kasey, Katelyn, Katherine, Kathleen, Kathryn, Katie, Kaylee, Kelley, Kellie, Kelly, Kelsey, Kerry, Krista, Kristen, Kristi, Kristin, Kristine, Kylie, Laura, Lauren, Laurie, Leigh, Lindsay, Lindsey, Lori, Lynn, Mackenzie, Madeline, Madison, Mallory, Maureen, Meagan, Megan, Meghan, Meredith, Misty, Molly, Paige, Rachael, Rebecca, Rebekah, Sara, Sarah, Savannah, Susan, Suzanne
Black female
Alfreda, Amari, Aniya, Aniyah, Aretha, Ashanti, Ayana, Ayanna, Chiquita, Dasia, Deasia, Deja, Demetria, Demetrice, Denisha, Domonique, Eboni, Ebony, Essence, Iesha, Imani, Jaleesa, Jalisa, Janiya, Kenisha, Kenya, Kenyatta, Kenyetta, Keosha, Keyona, Khadijah, Lakeisha, Lakesha, Lakeshia, Lakisha, Laquisha, Laquita, Lashanda, Lashawn, Lashonda, Latanya, Latasha, Latesha, Latisha, Latonia, Latonya, Latoria, Latosha, Latoya, Latrice, Mahogany, Marquita, Nakia, Nikia, Niya, Nyasia, Octavia, Precious, Quiana, Rashida, Sade, Shakira, Shalonda, Shameka, Shamika, Shaneka, Shanequa, Shanice, Shanika, Shaniqua, Shanita, Shaniya, Shante, Shaquana, Sharita, Sharonda, Shavon, Shawanda, Sherika, Sherita, Tameka, Tamia, Tamika, Tanesha, Tanika, Tanisha, Tarsha, Tawanda, Tawanna, Tenisha, Thomasina, Tierra, Tomeka, Tomika, Towanda, Toya, Tyesha, Unique, Willie, Zaria
A.2 Professions
We list the professions/fields used in our analysis:
Generated (Resumes)
Account Executive, Accountant, Administrative Assistant, Back-End Developer, Data Analyst, Data Engineer, Data Scientist, Firmware Engineer, Front-End Developer, Graphic Designer, Hardware Engineer, Legal Counsel, Marketing Manager, Mobile Developer, PR Specialist, Product Manager, Quality Assurance Engineer, Recruiter, Research Scientist, Supply Chain Manager, Technical Writer, UX Designer
Generated (Job Posts)
Account Executive, Data Analyst, Data Scientist, Firmware Engineer, Graphic Designer, Marketing Manager, Product Manager, Research Scientist, Supply Chain Manager, Technical Writer, UX Designer
Kaggle (Resumes)
Accountant, Advocate, Agriculture, Apparel, Arts, Automobile, Aviation, Banking, BPO, Business Development, Chef, Construction, Consultant, Designer, Digital Media, Engineering, Finance, Fitness, Healthcare, HR, Information Technology, Public Relations, Sales, Teacher
Kaggle (Job Posts)
Apparel, Aviation, Banking, Chef, Construction, Consultant, Finance, Fitness, Healthcare, IT, Teacher
A.3 Prompts
Generating Seed Resumes
Can you create a resume for the role [PROF] using the example below as a template? In addition, your generated resume should follow these instructions:
  • Use the resume below as a guide, but modify its content so that the generated resume is sufficiently different. Do not copy directly from the example resume.
  • Please tailor the education and skills to the role [PROF], and choose from a diverse set of schools and companies. Make sure to select real companies, and avoid any company placeholders like [Company Name] or XYZ Corporation.
  • Remove personal information from the resume, and do not generate any names, emails, websites, or addresses.
  • Do not include any placeholders for personal information (e.g., <PHONE_NUMBER>, <PERSON>, [Name], [Address], [Email], etc.).
  • The top of the resume should display [PROF].
  • Output a complete and detailed resume that follows a similar length and format to the example resume. Output only the resume, and do not include any text before or after it.
Adding Extracurriculars
You are a/an [RACE], [GENDER] professional. Can you add three sections to the resume below: (1) AWARDS, (2) MENTORSHIP AND VOLUNTEERING, and (3) CLUBS AND ORGANIZATIONS, that are reflective of your identity and background? If any of these sections (or equivalent sections) are already present, please modify them to better reflect the experiences of a/an [RACE], [GENDER] professional. Please output the full revised resume, which includes these additional sections. The other sections should remain the same, and please keep the overall format the same as well. Keep [NAME] at the top of the resume, and also add ‘——–’ at the top and bottom of the resume. Do not output anything else before or after the resume.
Generating Summaries
Generate a [LENGTH]-word summary of the resume above and its suitability for the role: [PROF]. Make sure the format is a paragraph and written in [POV] person.
A.4 Example of Extracurricular Information
We present generated extracurricular activities for Latoya Williams vs. Heather Williams (Product Manager), given the same initial resume. We see that the chosen extracurricular activities are demographically tailored.
Latoya Williams
Mentorship and Volunteering
  • Volunteer mentor at Black Girls Code, dedicated to empowering young Black women in exploring technological innovations and fostering an inclusive tech environment.
  • One-on-one mentorship program with aspiring product managers at Meta Platforms, offering guidance and support to enhance their professional growth.
Clubs and Organizations
  • Member of the Stanford University Black Students Association, fostering a supportive community and promoting cultural awareness on campus.
  • Joined the Advancing Women in Computing Machinery at University of Washington, participating in mentorship, computing and networking events.
Heather Williams
Mentorship and Volunteering
  • Volunteer Mentor, Girls Who Code - Guided and inspired high school girls interested in technology, encouraging them to pursue STEM careers.
  • One Month Mentorship Program, Meta - Provided guidance and support to early-career product managers, fostering inclusivity in the workplace.
Clubs and Organizations
  • Member, Stanford University Women in Business Society - Connected with like-minded professionals and promoted gender equality in the workplace.
  • Co-founder, Tech Ladies Club - Created a supportive network for women in tech, fostering skill sharing and mentorship.
A.5 Proxy Measures
We use the following measures as proxies for undesirable variation that could influence the decision of an HR staff reading the summary:
  • Reading ease is measured using Flesch Reading Ease score (Kincaid, 1975), with higher scores indicating greater ease. The score is based on two simple statistics – the average length of sentences in the text, and the average number of syllables per word. [^8] https://pypi.org/project/textstat/
  • Reading time is proportional to the number of characters in the text, with each character assigned a constant time to process. Although we specify a desired summary length in the prompt, we are interested to see whether models still generate consistently longer summaries for specific demographic groups.
  • Polarity quantifies the sentiment in text. We use Textblob’s implementation, [^9] https://pypi.org/project/textblob/ which returns scores closer to -1 for negative sentiment and scores closer to 1 for positive sentiment.
  • Subjectivity quantifies how much personal opinion vs. factual information is present in the text. Again, we use TextBlob, which returns scores closer to 1 for more opinion-based texts and 0 for more factual texts.
  • Regard captures whether a demographic group is positively or negatively perceived (Sheng et al., 2019). Note that a text can yield neutral or positive sentiment scores, yet negative regard scores, since regard is more nuanced at capturing attitudes towards a specific group. We utilize the regard classifier provided by Sheng et al. (2019).
(a) Quantification
(b) Focus
(c) Individual Impact
Figure 7: Human Annotation Results for 3 characteristics (quantification, focus, and individual impact).
A.6 Human Preferences
It is unclear whether the chosen measures for summarization (reading ease, reading time, polarity, subjectivity, and regard) capture meaningful differences in summaries. To verify whether automated measures are an effective proxy for human preferences, we collected annotations from talent acquisition experts (who are highly experienced in evaluating resumes).
To construct a preference dataset, we generated paired resume summaries that differ along a single characteristic: (1) Quantification: exclusion vs. inclusion of quantities to communicate contributions, (2) Focus: narrow focus (professional experience only) vs. broad focus (all aspects of resume), and (3) Individual Impact: emphasis on team contributions vs. individual impact. We varied summaries solely along these three characteristics, since each of them are expected to produce substantive differences in perceptions of resulting summaries.
We then asked experts [^10] We recruited 6 HR professionals to be annotators (US, Canada, and UK based), and conveyed that annotations would be used towards research on evaluating LLMs in hiring pipelines. We did not provide any monetary compensation to annotate the preferred summary in the pair (200 pairs annotated in total), and investigated whether experts displayed consistent preferences with respect to the characteristics being varied (quantification, focus, and individual impact). We gave the following instructions:
Overview: We would like to better understand the characteristics that contribute to good resume summaries. Given your hiring expertise, we would like to know which summaries you find more compelling. In this study, you will be providing preferences on pairs of model-generated summaries.
Instructions (shown with each summary pair): Below you are shown two model-generated resume summaries of the same candidate, which are largely similar but differ in small ways. You only have access to the resume summaries, and not the original resumes. Which resume summary below do you prefer?
We find that 4 out of 6 annotators favor the use of quantification, while 1 annotator prefer no quantification (Appendix Figure 7(a)). We see that 4 out of 6 annotators demonstrate a modest preference for focus, with the other 2 remaining neutral (Appendix Figure 7(b)). Additionally, 3 out of 6 annotators display a slight preference for individual impact, while 1 annotator displays a strong preference against it (Appendix Figure 7(c)). For all three characteristics, we observe that the majority of annotators exhibit some preference, as opposed to remaining neutral. Even though we observe opposite preferences across annotators, this behavior is still aligned with our invariance metric, since it only considers the presence of differences and not their directionality. Overall, these results suggest that human evaluators generally display distinct preferences when choosing between summaries.
Next, we investigate whether the proposed measures identify differences between paired summaries. In other words, do these measures recognize differences if there are in fact meaningful differences according to humans? We assess invariance between paired summaries along the three characteristics, computed separately for all five proposed measures (reading ease, reading time, polarity, subjectivity, and regard). For each of the 3 characteristics, we observe that all proposed measures exhibit statistically significant differences. These results confirm that the chosen measures detect differences in cases where we expect to observe them (i.e., based on results from human preferences).
(a) Top-5%
(b) Top-10%
Figure 8: Non-uniformity metric for top-5 and top-10% of retrieved Kaggle resumes. Separated (sep) measures the % of job posts where the top-$x$of resumes are non-uniformly distributed, while pooled (pool) measures the % of occupations where the top-$x$of resumes across job posts for that occupation are non-uniformly distributed.
A.7 Retrieval Fairness Metrics
We would like to emphasize that non-uniformity and exclusion are complementary rather than redundant metrics. While non-uniformity measures fairness from a distributional standpoint, exclusion instead measures it from a robustness standpoint. Intuitively, they answer different questions:
Non-uniformity
Let us consider demographically perturbed but otherwise equivalent resumes for four demographic groups: Black female, White female, Black male, and White male. Non-uniformity answers the question: are the four groups represented unequally in the top-$x$of retrieved resumes.
Exclusion
Let us consider the top-$n$ White male resumes for a given job post. Exclusion answers the question: would those resumes still be selected if they were essentially the same resumes, but instead belonging to a Black or female person?
As we see in Figures 2 and 3, the two metrics lead to different conclusions about the best retrieval model in terms of fairness (text-embedding-3-small for non-uniformity vs. embed-english-v3.0 for exclusion). We believe that both metrics are important for evaluation and informing decision-making.
(a) Gender, $n=5$
(b) Gender, $n=10$
(c) Gender, $n=100$
(d) Race, $n=5$
(e) Race, $n=10$
(f) Race, $n=100$
Figure 9: Directional differences in exclusion metric for retrieval (generated resumes) after applying name perturbations (i.e., separating based on perturbation direction). M→F perturbs male names to female names and F→M perturbs female names to male names, while W→B perturbs White names to Black names and B→W perturbs Black names to White names.
(a) Within-group Name, $n=5$
(b) Within-group Name, $n=10$
(c) Within-group Name, $n=100$
(d) Non-name, $n=5$
(e) Non-name, $n=10$
(f) Non-name, $n=100$
Figure 10: Exclusion metric for retrieval after performing non-demographic perturbations on generated resumes (i.e., within group name changes - top, and modifying spacing and adding typos - bottom).
(a) Gender, $n=5$
(b) Gender, $n=10$
(c) Gender, $n=100$
(d) Race, $n=5$
(e) Race, $n=10$
(f) Race, $n=100$
Figure 11: Directional differences in exclusion metric for retrieval (Kaggle resumes) after applying name perturbations (i.e., separating based on perturbation direction). M→F perturbs male names to female names and F→M perturbs female names to male names, while W→B perturbs White names to Black names and B→W perturbs Black names to White names.
(a) Within-group Name, $n=5$
(b) Within-group Name, $n=10$
(c) Within-group Name, $n=100$
(d) Non-name, $n=5$
(e) Non-name, $n=10$
(f) Non-name, $n=100$
Figure 12: Exclusion metric for retrieval after performing non-demographic perturbations on Kaggle resumes (i.e., within group name changes - top, and modifying spacing and adding typos - bottom).

### v2

@@
-Who Does the Giant Number Pile Like Best: Analyzing Fairness in Hiring Contexts
+Small Changes, Large Consequences: Analyzing the Allocational Fairness of LLMs in Hiring Contexts
 Abstract
-Large language models (LLMs) are increasingly being deployed in high-stakes applications like hiring, yet their potential for unfair decision-making and outcomes remains understudied, particularly in generative settings.
-In this work, we examine the fairness of LLM-based hiring systems through two real-world tasks: resume summarization and retrieval.
-By constructing a synthetic resume dataset and curating job postings, we investigate whether model behavior differs across demographic groups and is sensitive to demographic perturbations.
-Our findings reveal that race-based differences appear in approximately 10% of generated summaries, while gender-based differences occur in only 1%.
-In the retrieval setting, all evaluated models display non-uniform selection patterns across demographic groups and exhibit high sensitivity to both gender and race-based perturbations.
-Surprisingly, retrieval models demonstrate comparable sensitivity to non-demographic changes, suggesting that fairness issues may stem, in part, from general brittleness issues.
-Overall, our results indicate that LLM-based hiring systems, especially at the retrieval stage, can exhibit notable biases that lead to discriminatory outcomes in real-world contexts.
-Who Does the Giant Number Pile Like Best: Analyzing Fairness in Hiring Contexts
-Preethi Seshadri UC Irvine preethis@uci.edu                        Seraphina Goldfarb-Tarrant Cohere seraphina@cohere.com
+Large language models (LLMs) are increasingly being deployed in high-stakes applications like hiring, yet their potential for unfair decision-making remains understudied in generative and retrieval settings.
+In this work, we examine the allocational fairness of LLM-based hiring systems through two tasks that reflect actual HR usage: resume summarization and applicant ranking.
+By constructing a synthetic resume dataset with controlled perturbations and curating job postings, we investigate whether model behavior differs across demographic groups.
+Our findings reveal that generated summaries exhibit meaningful differences more frequently for race than for gender perturbations.
+Models also display non-uniform retrieval selection patterns across demographic groups and exhibit high ranking sensitivity to both gender and race perturbations.
+Surprisingly, retrieval models can show comparable sensitivity to both demographic and non-demographic changes, suggesting that fairness issues may stem from broader model brittleness.
+Overall, our results indicate that LLM-based hiring systems, especially in the retrieval stage, can exhibit notable biases that lead to discriminatory outcomes in real-world contexts.
+Small Changes, Large Consequences:
+Analyzing the Allocational Fairness of LLMs in Hiring Contexts
+Preethi Seshadri UC Irvine preethis@uci.edu          Hongyu Chen Cohere charlie@cohere.com          Sameer Singh UC Irvine sameer@uci.edu          Seraphina Goldfarb-Tarrant Cohere seraphina@cohere.com
 1 Introduction
-Large language models (LLMs) are increasingly being used for real-world, high-stakes domains, including medicine, law, and hiring.
-As LLMs are incorporated into critical decision-making processes, ensuring reliable and responsible usage is paramount.
-A key aspect of this effort includes anticipating and preventing specific risks and harms, such as identifying biases that could unfairly disadvantage certain demographic groups in hiring screening processes.
-Prior work focuses on two types of harms: representational harms (i.e., representing certain groups in a less favorable light than others, demeaning them, or failing to recognize their existence altogether) and allocational harms (i.e., allocating resources or opportunities or unfairly to different social groups) (Barocas et al., 2017; Blodgett et al., 2020).
-There is a substantial body of work that analyzes representational harms in LLMs (Zhao et al., 2018; Abid et al., 2021; Kirk et al., 2021; Cheng et al., 2023; Gadiraju et al., 2023), yet allocational harms—which are the primary harm at play in high-stakes situations—remain understudied.
-Given the growing use of LLMs in consequential applications, it is especially important to examine whether models produce unfair outcomes or performance gaps between different groups (allocational harms, also referred to as allocational fairness).
-Although there are a few studies on LLM harms in real-world settings (e.g., hiring) (Tamkin et al., 2023; An et al., 2024; Haim et al., 2024; Nghiem et al., 2024), there is minimal work closely simulating actual deployment scenarios.
-By using highly simplified setups and tasks, these prior works may fail to predict real-world outcomes and as a result provide an incomplete assessment of harms.
-Investigations into LLM harms should ensure ecological validity (Goldfarb-Tarrant et al., 2021; Cao et al., 2022); they should either be grounded in realistic scenarios and tasks that match how these systems are used in practice, or use a proxy that has been proven to be predictive of real world outcomes.
-Additionally, these existing studies have cast their investigations of allocational harms in LLMs as discrete classification tasks (such as binary accept/reject decisions) or quantitative predictions (such as determining salary levels).
-This makes the implementation and evaluation more straightforward, but does not match the real use of generative LLMs (Kelly, 2023; Yin et al., 2024).
-In contrast, there is limited work on allocational harms in generative settings (Wan et al., 2023), like summarization, since measuring how generated text might create disparities is more open-ended and complex than analyzing classification predictions.
-In this work, we examine whether LLMs are fair in real-world contexts for hiring.
-We choose hiring due to the growing adoption of LLMs in hiring tools (Sánchez-Monedero et al., 2020).
-Hiring is a high-stakes domain with clear societal risks – unfair systems may discriminate against already marginalized groups, risking further underrepresentation throughout the hiring pipeline (Dastin, 2018; Raghavan et al., 2020; Suresh and Guttag, 2021).
-We focus on two tasks that mirror how LLMs are integrated into hiring and recruitment workflows: resume summarization and resume retrieval with respect to a job post (Herman, 2024; Humanly, 2024).
-We then examine whether model behavior in these tasks is sensitive to gender and race demographic perturbations in resumes (e.g., name perturbations or tailored extracurricular content).
-Specifically, we ask the following questions:
-  • RQ1: Do model generated summaries differ meaningfully across demographic groups?
-  • RQ2: In a retrieval setting, do models disparately select resumes across demographic groups?
-  • RQ3: In a retrieval setting, how sensitive are model rankings to demographic and non-demographic perturbations in resumes?
-In this work, we: (1) create a synthetic resume dataset and curate a job post dataset, (2) propose metrics to measure fairness in generative and retrieval settings for two real-world hiring tasks, validated by an expert human preference study, and (3) perform a study investigating the fairness of an automated LLM hiring system based on real world usage, for 7 models (3 generative, 4 retrieval).
-[^1] The data and code can be found at https://github.com/preethisesh/hiring_fairness.
-In the summarization setting, on average, we observe meaningful differences between summaries generated for different racial groups nearly 10% of the time, as opposed to only 1% for gender (RQ1).
-On the other hand, for retrieval, we find that all models are prone to non-uniformly selecting resumes across demographic groups (RQ2).
-In addition, all models produce rankings that are highly sensitive to both name and race demographic perturbations (RQ3).
-We also find that model sensitivity to non-demographic perturbations is notable and sometimes even on par with demographic perturbations (RQ3), suggesting that fairness issues may largely be due to model brittleness as a whole, rather than encoded biases only.
-Overall, our results demonstrate that an LLM hiring system with automated resume retrieval and summarization exhibits considerable bias, primarily stemming from the initial retrieval stage.
+Large language models (LLMs) are increasingly being adopted in real-world, high-stakes domains such as hiring (Boston Consulting Group, 2025), where they assist HR teams with tasks like resume screening and candidate matching.
+As LLMs are incorporated into critical decision-making processes, ensuring fair and responsible deployment is essential, especially when the outcomes can profoundly impact individuals’ career prospects (Dastin, 2018; Raghavan et al., 2020; Sánchez-Monedero et al., 2020; Suresh and Guttag, 2021).
+A key aspect of developing responsible LLM systems includes anticipating and preventing specific risks and harms, such as allocational harms (i.e., allocating resources or opportunities unfairly to different social groups, also called allocational fairness) (Barocas et al., 2017; Blodgett et al., 2020).
+This is especially important in automated hiring pipelines, since models may produce unfair outcomes and reinforce systemic inequalities.
+While there is a substantial body of work that analyzes representational harms (i.e., representing certain social groups negatively, demeaning them, or erasing their existence) in LLMs (Zhao et al., 2018; Abid et al., 2021; Kirk et al., 2021; Cheng et al., 2023; Gadiraju et al., 2023), allocational harms—which are the primary harm at play in high-stakes situations—remain understudied beyond discriminative systems.
+The few studies that evaluate allocational harms of LLMs (Tamkin et al., 2023; An et al., 2024; Haim et al., 2024; Nghiem et al., 2024) have primarily cast their investigations as discrete classification tasks (e.g., yes/no decisions) or quantitative predictions (e.g., determining salary levels), which do not capture how LLMs are deployed in applications like hiring (Kelly, 2023).
+As a result, these highly simplified setups may inadequately predict real-world outcomes and assess harms.
+Investigations of LLM harms must ensure ecological validity (Blodgett et al., 2021; Goldfarb-Tarrant et al., 2021; Cao et al., 2022); they should be grounded in realistic scenarios and tasks that match how these systems are used in practice, or use a proxy that is predictive of real world outcomes.
+Yet there is limited work on allocational harms in generative settings without adding a simplification layer, with Wan et al.
+(2023) being a notable exception, since measuring how generated text might create disparities is more open-ended and complex than analyzing classification predictions.
+In this work, we examine whether LLMs behave fairly in real-world hiring contexts.
+We focus on two critical tasks that mirror how LLMs are integrated into hiring workflows (Herman, 2024; Humanly, 2024): (1) ranking candidates with respect to a job posting and (2) summarizing resumes, as illustrated in Figure 1.
+These tasks represent key stages where automation can influence which candidates are surfaced and considered for a role.
+To evaluate fairness, we examine whether models are sensitive to gender and race perturbations in resumes.
+We investigate the following questions:
+  • RQ1: Do generated summaries differ meaningfully across demographic groups? [^1] We study summarization first, since it is less explored from an allocational harms perspective.
+  • RQ2: Do models disparately select resumes across demographic groups?
+  • RQ3: How sensitive are model rankings to demographic and non-demographic perturbations in resumes?
+Figure 1: We investigate the fairness of an LLM hiring pipeline with a retrieval stage (ranks/filters the top-$n$ candidates with respect to a job post) and a summarization stage (generates resume summaries for filtered candidates).
+We assess fairness at each stage separately.11footnotemark: 1
+To this end, we: (1) construct a new benchmark consisting of a synthetic resume dataset with controlled demographic perturbations (varying names and extracurricular content) and curated job postings, (2) design an evaluation framework with fairness metrics tailored to both generative and retrieval settings, validated by an expert human preference study, (3) conduct a comprehensive fairness analysis of 10 large language models (6 generative, 4 retrieval) based on real-world hiring tasks.
+We will make all data and code publicly available.
+Our results demonstrate that an LLM hiring system with automated resume retrieval and summarization exhibits considerable bias, primarily stemming from the retrieval stage.
+In the summarization setting, we observe meaningful differences in generated summaries up to 20% of the time between racial groups, compared to 3% for gender (RQ1).
+For retrieval, models non-uniformly select resumes across demographic groups up to 55% of the time (RQ2), and produce rankings that are highly sensitive to gender and race, with up to 74% of candidates being filtered out after demographic perturbation (RQ3).
+We also find that models exhibit high sensitivity to non-demographic changes, sometimes on par with demographic changes (RQ3), suggesting that fairness issues can stem from general model brittleness rather than demographic bias alone.
+Overall, our analysis reveals that even seemingly minor changes can lead to considerable disparities, raising concerns about the fairness and robustness of LLMs in hiring.
 2 Methodology
-To study fairness in hiring, we consider an LLM-based pipeline with two components: resume retrieval with respect to a job post (using an embedding model), and resume summarization (using an LLM).
-[^2] We analyze both components because they reflect real-world usage of automation to streamline hiring processes.
-We focus on summarization first because it is more neglected in research, though in a pipeline it would come after retrieval (as summarization would be of retrieved resumes).
-We investigate the impact of perturbing the demographic group of a resume on each component of the pipeline and look for: (1) systematic differences in generated resume summaries and (2) changes in similarity of a resume to a job posting, and as a result its ranking in a resume set.
+To study fairness in hiring, we consider an LLM-based pipeline with two components: resume retrieval with respect to a job post (using an embedding model) and resume summarization (using an LLM).
+This pipeline is informed by interviews with several corporations that actively deploy LLMs for hiring; [^2] We cannot share details due to non-disclosure agreements.
+both components reflect real-world usage of automation to streamline hiring processes.
+We focus on summarization first because it is more neglected in research, though in a pipeline it would come after retrieval, as shown in Figure 1 (as summarization would be of retrieved resumes).
+We propose two metrics, invariance violations (summarization) and exclusion (retrieval), to investigate allocational fairness in hiring.
+Specifically, these metrics quantify: (1) systematic differences in generated resume summaries [^3] Since recruiters may rely on summaries rather than full resumes, meaningful differences in summaries could directly affect hiring outcomes, leading to allocational harms.
+and (2) changes in the similarity of a resume to a job posting, and as a result its ranking in a resume set.
+We additionally benchmark the distribution metric of Wilson and Caliskan (2024) to study fairness in hiring, but note that their approach does not directly capture how perturbing a resume impacts resume screening outcomes, and includes only retrieval and not generative settings.
 Let $D$ represent a set of resumes, where each resume $d\in D$ has demographic label $l(d)$.
-For each resume, let $d^{\prime}$ represent any perturbed version of $d$ where $l(d^{\prime})\neq l(d)$.
-For each $d$, there can be several perturbed versions $d^{\prime}$ along various demographic dimensions (e.g., gender or race, to name a few).
+We denote $d^{\prime}$ as any perturbed version of $d$ where $l(d^{\prime})\neq l(d)$.
+For each $d$, there can be several demographically perturbed versions $d^{\prime}$ (e.g., gender or race perturbations).
 The resume content remains largely unchanged except for the demographic perturbation.
 2.1 Summarization
-Perturbed resumes are by design highly similar to original resumes, so we expect generated summaries for original and perturbed resumes to also maintain a high degree of similarity.
-In other words, we are testing for invariance, since we expect the output to change minimally after perturbing the input (Ribeiro et al., 2020).
-In order to assess invariance, we need a way to measure whether the original and perturbed summaries differ meaningfully.
-For cost and scalability reasons, we rely on an automated approach—human preferences are expensive and cannot be collected frequently, especially if these preferences are sourced from HR staff.
-Additionally, collecting human preferences for a specific dataset does not extend to other analyses with new datasets.
-Automated measures can evaluate whether specific properties of generated summaries differ, where differences in these properties would be undesirable in practice.
-For instance, summaries for a specific demographic group should not be written more positively than summaries for other demographic groups; this could lead to disparity in hiring outcomes.
-We use the following measures as proxies for undesirable variation that could influence the decision of an HR staff reading the summary: reading ease, reading time, polarity, subjectivity, and regard (Appendix A.5).
+Perturbed resumes are by design highly similar to original resumes, so we expect generated summaries for original and perturbed resumes to also maintain high similarity.
+In other words, we are testing for invariance; we expect the output to change minimally after perturbing the input (Ribeiro et al., 2020).
+To assess invariance, we need a way to measure whether the original and perturbed summaries differ meaningfully in the context of hiring.
+For cost and scalability reasons, we rely on an automated approach—human preferences are expensive and cannot be collected quickly enough to be used in model development, especially if these preferences are sourced from HR staff.
+Automated measures can evaluate whether specific properties of generated summaries differ, in a way that could affect a human reader’s opinion.
+For instance, summaries for a specific demographic group should not be written more positively than summaries for other demographic groups; this could lead to disparities in hiring outcomes.
+We use the following measures as proxies for undesirable variation that could influence the decision of HR staff reading the summary: reading ease, reading time, polarity, subjectivity, and regard (Appendix A.8).
 We verify that these measures capture meaningful differences in practice by conducting a preference task annotated by HR staff.
@@
 Fairness Metric
-We compute an invariance metric that aggregates the proposed measures for both original and perturbed resume summaries.
-For each measure, we do a paired t-test between scores for original and perturbed summaries, repeated for each demographic perturbation.
-We then calculate how often the null hypothesis (that the mean difference between paired summaries is zero) is rejected, i.e., how often invariance is violated.
-We choose a significance level of $\alpha=0.05$, and apply Bonferroni correction to account for Type 1 errors with multiple comparisons (Bland and Altman, 1995).
+For each measure, we perform a paired t-test between scores for original and perturbed summaries.
+We then calculate how often the null hypothesis (that the mean difference between paired summaries is 0) is rejected, i.e., how often invariance is violated.
+We choose a significance level of $\alpha=0.05$, and apply Benjamini-Hochberg correction to account for Type 1 errors with multiple comparisons (Benjamini and Hochberg, 1995).
 […]
 2.2 Retrieval
-In contrast to summarization, which relies on only the resume, retrieval utilizes both a query (i.e., job posting) and candidates to select (i.e., resumes).
+In contrast to summarization, which relies only on the resume, retrieval uses both a query (job posting) and candidates to select (resumes).
 The top-$n$ of the resulting resumes sorted by similarity then make it to the next stage of the pipeline.
-Similar to the previous setting, we assume that demographically perturbing a resume should have no impact on its relevance to a job posting.
-Given an embedding model $\mathcal{M}$ and a similarity measure $sim$, we compute the similarity between embeddings for resume $d$ and a job posting $p$.
-In practice, we use cosine similarity (also used in Wilson and Caliskan (2024)).
-We define the set $S(p)$, which represents the set of similarity values between each resume $d\in D$ and job posting $p$:
-$S(p)=\{sim(\mathcal{M}(d),\mathcal{M}(p))\mid d\in D\}$
-We transform the similarity values between each resume and job posting, $s_{i}$, into a rank such that lower ranks indicate higher similarity (note: the highest similarity value receives rank 1):
-$\text{rank}_{p}(s_{i})=|\{s_{j}\in S(p)\mid s_{j}>s_{i}\}|+1$
-Let $D_{n}(p)$ represent the set of top-$n$ resumes from $D$, which are the resumes with the $n$ lowest ranks (i.e., highest similarities) for job posting $p$:
-$D_{n}(p)=\{d_{i}\in D\mid\text{rank}(s_{i})\leq n\}$
+We assume that demographically perturbing a resume should have minimal impact on its relevance to a job posting.
+Given an embedding model $\mathcal{M}$ and a similarity measure simsim, we compute the similarity between embeddings for resume $d$ and job posting $p$.
+In practice, we use cosine similarity following Wilson and Caliskan (2024).
+We define the set $S(p)$, which represents the set of similarity values between each resume $d\in D$ and a given job posting $p$: $S(p)=\{sim(\mathcal{M}(d),\mathcal{M}(p))\mid d\in D\}$.
+We transform the similarity values between each resume and job posting, $s_{i}$, into a rank such that lower ranks indicate higher similarity: $\text{rank}_{p}(s_{i})=|\{s_{j}\in S(p)\mid s_{j}>s_{i}\}|+1$.
+Let $D_{n}(p)$ represent the set of top-$n$ resumes from $D$, which are the resumes with the $n$ lowest ranks (i.e., highest similarities) for job posting $p$: $D_{n}(p)=\{d_{i}\in D\mid\text{rank}(s_{i})\leq n\}$.
 Fairness Metrics
 We compute two fairness metrics for retrieval, non-uniformity, proposed by Wilson and Caliskan (2024), and exclusion, which we introduce below.
-Further intuition about these metrics are provided in Appendix A.7.
+[^4] Further intuition and differentiation of these metrics are provided in Appendix A.11.
 Non-uniformity
-evaluates whether the top-$x$(where $x$ is a percentage instead of a fixed number $n$) of resumes are uniformly distributed with respect to the demographic groups being considered.
-To compute non-uniformity, we consider the set of top-$x$of resumes retrieved from the combined pool of all demographically perturbed versions, which we denote as $D_{x}^{\prime}(p)$.
-Following the approach from Wilson and Caliskan (2024), we perform a chi-squared goodness-of-fit test to determine whether the demographic makeup of $D_{x}^{\prime}(p)$ deviates from the uniform distribution.
+assesses whether the top resumes are uniformly distributed across demographic groups.
+First, the set of top-xx% of resumes ($x$ being a percentage rather than a fixed number $n$) is retrieved from the combined pool of all demographically perturbed versions, denoted as $D_{x}^{\prime}(p)$.
+A chi-squared goodness-of-fit test is then used to check if the demographic composition of $D_{x}^{\prime}(p)$ deviates from the uniform distribution.
 Exclusion
@@
 Ideally, $\mathcal{M}$ should be robust to demographic perturbations, yielding nearly identical similarity scores and rankings for both resume $d$ and demographically perturbed version $d^{\prime}$.
-Exclusion directly tests this behavior by measuring the extent to which the set of top-$n$ resumes differs after perturbation.
+Exclusion directly assesses allocational fairness by measuring how much the set of top-$n$ resumes differs after perturbation.
 $\text{exclusion}_{n}(p)=\frac{|{\text{rank}_{p}(d^{\prime})>n\mid d\in D_{n}(p)}|}{|D_{n}(p)|}$
 3 Experimental Setup
-In this section, we provide an overview of the datasets, perturbations, and models used in our evaluation.
+In this section, we describe the data, perturbations, and models used for our evaluation.
 3.1 Data
 Resumes
-Resumes are sourced through social media platforms including LinkedIn, Slack, and X.
+Resumes are sourced through social media platforms (LinkedIn, Slack, X).
 Given the authors’ professional circles, the sample skews heavily toward tech and academic professionals.
-For privacy reasons, we anonymize resumes using Presidio anonymize, [^3] https://microsoft.github.io/presidio/anonymizer/ which masks PII entities.
-To further mitigate privacy concerns and enable us to release the dataset, we use the collected resumes as examples to generate synthetic resumes, instead of using original resumes in our analysis.
-We generate 525 resumes across 21 professions using Cohere’s Command-R model (Cohere, 2024).
-All synthetic resumes are free of explicit demographic information, until we add it as part of experimentation.
-In addition to our curated resume set, we use a publicly available resume dataset from Kaggle (Bhawal, 2021).
-[^4] Has a CC0: Public Domain License These resumes differ in two notable ways: (1) the Kaggle resumes are less structured and formatted than synthetically generated resumes, and (2) they include a more diverse set of fields (e.g., apparel, construction, fitness, etc.).
+For privacy reasons, we anonymize resumes using Presidio to mask PII entities (Microsoft, ).
+To further mitigate privacy concerns and enable dataset release, we use the collected resumes as examples to generate synthetic resumes.
+We generate 525 resumes across 22 professions using Cohere’s Command-R model (Cohere, 2024).
+All synthetic resumes are free of explicit demographic information, until added during experimentation.
+In addition, we use a publicly available resume dataset from Kaggle (Bhawal, 2021) to increase coverage and generalization.
+These resumes differ in two notable ways: (1) they are less structured and formatted than generated ones, and (2) they include a more diverse set of fields (e.g., construction, fitness, etc.).
 We sample 1175 Kaggle resumes across 24 fields.
+More details about dataset curation and statistics are provided in Appendix A.4 - A.6.
 Job Posts
-Our resume dataset consists of two types: realistic synthetic resumes generated for specific roles (e.g., data analyst) and actual resumes from Kaggle labeled with broader field categories (e.g., construction).
-For each profession/field (we choose 11 from generated resumes and 11 from Kaggle resumes), we carefully select 7 detailed LinkedIn job postings that clearly describe the role and outline qualifications.
-For Kaggle resumes, we match postings to roles within each broader field (e.g., choosing “construction manager” postings for the construction field).
-In total, we collect 154 job postings across different professions.
+Our resume dataset consists of two types: synthetic resumes generated for specific roles (e.g., data analyst) and actual resumes labeled with broader field categories (e.g., construction).
+For each profession/field (we choose 11 each from generated and Kaggle resumes), we carefully select 7 detailed LinkedIn job postings, resulting in 154 job postings.
 3.2 Demographic Perturbations
 We use names as a proxy for gender and racial information.
-Initially, all resumes are free of names.
-We then add names using the curated set of names from Yin et al.
+All resumes are initially free of names; we add them using the curated set from Yin et al.
 (2024).
@@
 We consider four demographic groups, each with 100 unique names: Black female (FB), White female (FW), Black male (MB), and White male (MW).
-Following Wilson and Caliskan (2024), we use the demographically neutral “Williams” as the last name for all groups and only vary first name.
-In actual resumes, demographic information can be encoded in more than just the name.
-Therefore, for synthetically generated resumes, we perform an additional augmentation step that adds extracurricular information to the resumes (similar to Glazko et al.
-(2024)).
-We use Command-R to augment resumes by including awards, clubs and leadership, and mentorship and volunteering experiences that are reflective of the individual’s background and identity (more details in Appendix A.3 and A.4).
-In addition to explicit gender and race information from names, extracurricular information can reinforce demographic signal by providing further explicit and implicit cues.
-(a) GPT-4o
-(b) Command-R
-(c) Mixtral 8x7b
-Figure 1: Summarization Results: Percentage of t-tests for which we reject null hypothesis (i.e., invariance is violated), separated by completion model and perturbation.
+Following Wilson and Caliskan (2024), we only vary the first name and fix “Williams” as the last name for all groups.
+In actual resumes, demographic information can be encoded in more than just names.
+Therefore we perform an additional augmentation step that adds extracurricular information using Command-R [^6] Awards, clubs and leadership, and mentorship and volunteering experiences that are reflective of the individual’s background and identity (see Appendix A.6 and A.7).
+to the resumes (similar to Glazko et al.
+(2024)) Adding this information can reinforce demographic signal by providing both explicit and implicit cues.
 3.3 Non-Demographic Perturbations
-We do two non-demographic perturbation experiments for retrieval to assess baseline sensitivity of embedding models that is not due to demographics.
-Within Group Name Perturbations
-As a baseline comparison to performing name perturbations between different demographic groups (e.g., White female $\rightarrow$ Black Female), we assess whether models are sensitive to same group demographic perturbations (e.g., White female $\rightarrow$ White Female).
-By doing so, we disentangle how much bias is due demographic changes vs.
-overall model sensitivity to name changes.
+We conduct two non-demographic perturbation experiments for retrieval to assess the baseline sensitivity of embedding models that is not due to demographics.
+Within-Group Name Perturbations
+As a baseline comparison to performing name perturbations between different demographic groups (e.g., White female $\rightarrow$ Black Female), we assess whether models are sensitive to within-group demographic perturbations (e.g., White female $\rightarrow$ White Female).
+By doing so, we disentangle how much bias is due to demographics vs.
+model sensitivity to name changes.
 To control for the effects of frequency (Ethayarajh et al., 2019), we bin names in each demographic group according to their frequency in the Pile dataset (Gao et al., 2020), and match names based on the bin.
-Therefore, name perturbations should result in a similarly occurring name within the same demographic group.
+[^7] We use the What’s in My Big Data tool (Elazar et al., 2023) to obtain frequencies.
 Non-Name Perturbations
-We also assess whether model rankings are sensitive to non-name perturbations.
-This allows us to examine whether models lack robustness more broadly, or mainly exhibit brittleness with respect to name changes.
-We test two perturbation types: (1) random character swapping, which does not impact readability or comprehension in a meaningful way and (2) replacing new lines in the resume with a single space instead, which targets formatting without modifying content.
-For character swapping, we choose 10 random characters in the resume and swap based on neighboring keys to simulate typos.
+We assess whether model rankings are sensitive to non-name perturbations.
+This allows us to examine whether models lack robustness more broadly.
+We test two perturbation types: (1) random character swapping, which does not impact readability or comprehension in a meaningful way [^8] We choose 10 random characters in the resume and swap with neighboring keys to simulate typos.
+and (2) replacing new lines in the resume with a single space instead, which targets formatting without modifying content.
+(a) Gender
+(b) Race
+Figure 2: Summarization Results: Invariance violations for generated summaries, separated by completion model and perturbation type.
+Results are shown across 5 runs.
+Left 3 models are considered "smaller" models, right 3 models are considered "larger" models.
 3.4 Models
 Summarization
-We generate summaries using a combination of closed and open state-of-the-art language models: GPT-4o (OpenAI, 2024), Command-R (Cohere, 2024), and Mixtral 8x7B (Jiang et al., 2024).
-For summary instructions, we vary the generation length (100 and 200 words) as well as the point of view (first and third person) specified in the prompt.
+We generate summaries using closed and open state-of-the-art LLMs: GPT-4o (OpenAI, 2024), Command-R (Cohere, 2024), Mixtral 8x7B and Mistral Large (Jiang et al., 2024), and Llama 3.1 8B and Llama 3.3 70B (Meta AI, 2024).
+For summary instructions, we vary the generation length (100, 200 words) as well as the point of view (first, third person) specified in the prompt.
 We generate summaries with temperatures of 0.0 and 0.3.
+To account for stochasticity in generations, we generate each summary five times.
 Retrieval
-For retrieval, we select four popular dense embedding models used in retrieval augmented generation (RAG) systems (Lewis et al., 2020): OpenAI’s text-embedding-3-small and text-embedding-3-large models, Cohere’s embed-english-v3.0, and Mistral’s mistral-embed.
-(a) Top-5%
-(b) Top-10%
-Figure 2: Non-uniformity metric for top-5 and top-10% of retrieved resumes.
-Separated (sep) measures the % of individual job posts where the top-$x$of resumes are non-uniformly distributed, while pooled (pool) measures the % of occupations where the top-$x$of resumes across job posts for that occupation are non-uniformly distributed.
+For retrieval, we select four popular dense embedding models used in retrieval augmented generation (RAG) systems (Lewis et al., 2020): OpenAI’s text-embedding-3-small and text-embedding-3-large, Cohere’s embed-english-v3.0, and Mistral’s mistral-embed.
 4 Results
@@
 4.1 Summarization
-We analyze whether generated summaries differ meaningfully across demographic groups (RQ1) by examining invariance violations, i.e., the percentage of t-tests that yield significant differences in our automated measures.
-We measure across summaries with different characteristics (length, point of view, and temperature).
-Figure 1 displays results for generated resumes, grouped by completion model and demographic perturbation.
-Mixtral 8x7B exhibits the highest average invariance violations (8.59%), followed by GPT-4o (3.91%) and Command-R (2.74%).
-All models violate invariance much more for resumes that differ by race vs.
-gender.
-For example, the average invariance violations for race perturbations is 15.62% for Mixtral 8x7B, as compared to only 1.56% for gender.
-In summary, we observe that models exhibit some but not considerable discrepancies between generated summaries for different demographic groups, with different models having potentially large differences in sensitivity.
+We analyze whether generated summaries differ meaningfully when applying gender and race perturbations (RQ1) by examining invariance violations, i.e., the percentage of t-tests that yield significant differences in our automated measures.
+We measure violations separately for summaries with different characteristics (length, point of view, and temperature).
+Figure 2 displays results grouped by completion model and perturbation type.
+All models violate invariance much more for resumes that differ by race as opposed to gender.
+In fact, gender invariance violations are zero or near zero for all models.
+In contrast, all models except Command-R exhibit invariance violations with respect to race, with Mixtral 8x7B exhibiting violations 16.76% of the time on average.
+Our results also provide some indication that smaller models are more susceptible to violations.
+In summary, we observe that models exhibit some but not considerable discrepancies between generated summaries for different demographic groups, with minimal differences for gender perturbations.
 4.2 Retrieval
-We now shift to a retrieval setting and ask: do models exhibit fairness issues in resume retrieval?
-Our analysis tackles this question from both a distributional (RQ2) and robustness (RQ3) perspective.
+Moving to retrieval, we ask: Do models exhibit fairness issues in selecting resumes? Our analysis tackles this question from distributional (RQ2) and robustness (RQ3) perspectives.
 4.2.1 Non-Uniformity
 Do models disparately select resumes across demographic groups?
-We compute non-uniformity (i.e., how often top retrieved resumes have non-uniform demographic distributions).
-Non-uniformity differs considerably across models, choice of top-$x$ percent, and whether or not resumes are pooled across occupations (Figure 2).
+To answer this question, we compute non-uniformity (i.e., how often top retrieved resumes have non-uniform demographic distributions).
+All models disparately retrieve resumes across demographic groups, consistent with the findings of Wilson and Caliskan (2024).
+That being said, non-uniformity differs considerably across models, choice of top-$x$ percent, and pooling of resumes across occupations (Figure 3).
+We observe that embed-english-v3.0 exhibits the highest non-uniformity on average, with 6.90% of job posts and 45.45% of occupations having non-uniformly distributed resumes.
+Increasing top-$x$ from 5% to 10% and pooling resumes both yield higher non-uniformity across all models.
+In particular, pooling resumes by occupation can produce massive changes; on average across models, non-uniformity goes from 3.66% $\rightarrow$ 30.68%.
+This reflects sensitivity in the metric itself more than a change in the shape of the distribution.
+[^9] Increasing top-$x$ and pooling both increase sample size, which can lead to rejecting the null hypothesis in cases where the null hypothesis previously failed to be rejected.
+(c) Top-5%
+(d) Top-10%
+Figure 3: Non-uniformity metric for top-5 and top-10% of retrieved resumes.
+Separated (sep) measures non-uniformity at a job post level, while pooled (pool) measures it at an occupation level by pooling results across job posts for a given occupation.
 (a) $n=5$
@@
 Lower values indicate models are less sensitive to demographic perturbations.
-We observe that embed-english-v3.0 exhibits the highest non-uniformity on average, where 6.90% of job posts and 45.45% of occupations have non-uniformly distributed resumes.
-We see that increasing top-$x$ from 5% to 10% and pooling resumes both yield higher non-uniformity across all models.
-In particular, pooling resumes across occupations can produce massive changes – for example, we see that on average across models, non-uniformity goes from 3.66% $\rightarrow$ 30.68% after pooling, resulting in a nearly 750% increase.
-However, this reflects a sensitivity in the metric itself more than a change in the shape of the distribution.
-Increasing top-$x$ and pooling both increase sample size, which can lead to rejecting the null hypothesis that the distribution is uniform in cases where the null hypothesis previously failed to be rejected.
-Different models demonstrate distinct patterns of bias.
-For example, when considering the highest represented groups in the top-10% of resumes with respect to each job post, White females are the top group in 37 job posts (48.05%), while Black males are the top group in only 3 job posts (3.90%) for embed-english-v3.0.
-In contrast, for mistral-embed, White males are the top group in 56 job posts (72.73%), while White females are the top group in only 4 job posts (5.19%).
-We cannot determine why such differences exist without access to datasets or training details, but it is notable and somewhat surprising that models do not uniformly prefer one demographic group over the other.
+Different models show distinct patterns of bias: the non-uniformity privileges different demographic groups.
+For example, in the top-10% of resumes from embed-english-v3.0, White females are the top group 48.05% of the time, compared to 3.90% for Black males.
+In contrast, for mistral-embed, White males are the top group 72.73% of the time, compared to 5.19% for White females.
+Reasons for these differences are unclear without access to dataset and training details but it is notable and surprising that models do not consistently favor the same demographic group.
 4.2.2 Exclusion
 How sensitive are models to gender and race perturbations?
-We compute exclusion (i.e., how often top retrieved resumes are excluded from the set of top-$n$ resumes after demographic perturbations).
-We find that all models display notable sensitivity to gender and race name perturbations, as shown in Figure 3.
-When considering top-5 resumes, we find that models tend to exclude perturbed resumes nearly half the time (45.75% on average).
+We compute our proposed metric, exclusion (i.e., how often top retrieved resumes are excluded from the set of top-$n$ resumes after perturbations), and find that all models display notable sensitivity to gender and race name perturbations (Figure 4).
+When considering the top-5 resumes, we find that models tend to exclude perturbed resumes nearly half the time (45.75% on average).
 Across both gender and race name perturbations, and different $n$ values, text-embedding-3-small and text-embedding-3-large have the highest exclusion, while embed-english-v3.0 consistently has the lowest exclusion.
 As expected, exclusion lowers as $n$ increases, since larger $n$ values are less restrictive and consider a larger set of retrieved resumes.
-That being said, exclusion for $n=100$ is still considerable, as all models have exclusion > 12%.
+That being said, exclusion for $n=100$ is still considerable, as all models have exclusion >> 12%.
 [^10] In practice we expect $n$ to be low for filtering candidates.
-For exclusion, models have similar sensitivity to both gender and race perturbations.
-Overall, average exclusion for gender is 31.78% on generated resumes (25.47% on Kaggle resumes) vs.
-31.66% on generated resumes (25.40% on Kaggle resumes) for race.
-[^7] Kaggle resumes exhibit similar patterns but lower exclusion magnitude than generated resumes.
-This is likely because generated resumes are tech-focused and closer in content.
-These results contrast our summarization findings, where models show greater invariance violations on race vs.
-gender perturbations.
-Our analysis reveals that the set of top retrieved resumes with respect to a job posting is highly brittle, as merely altering the demographic on resumes (through names only) often results in otherwise identical resumes dropping out of the top-$n$ results.
 (a) Gender, $n=5$
@@
 M→F perturbs male to female names and F→M perturbs female to male names, while W→B perturbs White to Black names and B→W perturbs Black to White names.
-Does model sensitivity to perturbations differ based on the direction of perturbation?
-We can also partition the results based on direction of perturbation (Figure 4).
-We find that the magnitude of difference exceeds a threshold of 5% (e.g., the absolute difference in exclusion for M $\rightarrow$ F vs.
-F $\rightarrow$ M or W $\rightarrow$ B vs.
-B $\rightarrow$ W for the same model, $n$, and resume type is $>5$) in 26/48 cases across generated and Kaggle resumes, indicating models often exhibit higher sensitivity to one direction of perturbation over the other.
-In particular, the gender directional difference is notable for mistral-embed, going from 63.28% for M $\rightarrow$ F to 27.93% for F $\rightarrow$ M, for generated resumes with $n=5$ (Figure 4).
+In contrast to our summarization findings, where models show greater invariance violations for race vs.
+gender perturbations, models have similar sensitivity to gender vs.
+race perturbations for exclusion.
+Overall, average exclusion for gender is 31.78% on generated resumes (25.47% on Kaggle resumes) vs.
+31.66% on generated resumes (25.40% on Kaggle resumes) for race.
+[^11] Kaggle resumes exhibit similar patterns to generated resumes, but are lower in exclusion magnitude.
+This is likely because generated resumes are tech-focused and more overlapping in content.
+Our analysis reveals that the set of top retrieved resumes with respect to a job posting is highly brittle, as merely altering the demographic with names often results in otherwise identical resumes dropping out of the top-$n$ results.
 (a) Gender, $n=5$
 (b) Race, $n=5$
-Figure 5: Deltas (differences) in exclusion metric for retrieval after performing demographic perturbations with names as well as extracurricular information as compared to only names, separated by direction.
+Figure 6: Deltas (differences) in exclusion metric for retrieval after performing demographic perturbations with names + extracurricular information vs.
+names only.
 As expected, adding extracurricular information increases sensitivity to perturbations.
-We also observe that models exhibit opposite directional trends for gender and race in Figure 4.
-For gender, all models except mistral-embed are more sensitive when perturbing female names (marginalized) to male names (non-marginalized); for race, all models except mistral-embed are more sensitive when perturbing White names (non-marginalized) to Black names (marginalized).
-These results highlight an asymmetry in how models handle different types of demographic changes.
+Does model sensitivity to perturbations differ based on the direction of perturbation?
+We partition the results based on the perturbation direction (Figure 5), and find that models often exhibit higher sensitivity to one direction of perturbation over the other.
+In particular, the gender directional difference is notable for mistral-embed, going from 63.28% for M $\rightarrow$ F to 27.93% for F $\rightarrow$ M, for generated resumes with $n=5$.
+We also observe that models exhibit opposite directional trends for gender and race.
+For gender, all models except mistral-embed are more sensitive when perturbing female names (marginalized) to male names (non-marginalized).
+On the other hand, for race, all models except mistral-embed are more sensitive when perturbing White names (non-marginalized) to Black names (marginalized).
+These results highlight an asymmetry in how models handle various demographic changes.
 Are models more sensitive when perturbing both names and extracurricular information, as opposed to names only?
-Figure 5 shows that, as expected, models tend to be more sensitive when perturbing extracurricular information (e.g., awards, clubs, volunteering, etc.) in addition to names.
-On average, across models and $n$ values, we observe the following increases in exclusion: 9.35% for M $\rightarrow$ F, 8.06% for F $\rightarrow$ M, 16.41% for W $\rightarrow$ B, and 2.90% for B $\rightarrow$ W.
+Figure 6 shows that models tend to be more sensitive when perturbing extracurricular information in addition to names.
+On average, we observe the following increases in exclusion: 9.35% for M $\rightarrow$ F, 8.06% for F $\rightarrow$ M, 16.41% for W $\rightarrow$ B, and 2.90% for B $\rightarrow$ W.
 For gender, adding extracurricular information results in comparable increases in exclusion for both directions.
-In contrast, adding extracurricular information for race results in highly asymmetric increases in exclusion.
+In contrast, adding extracurricular information for race results in highly asymmetric increases.
 W $\rightarrow$ B averages more than 5x the increase of B $\rightarrow$ W changes.
@@
 This finding is notable given that prior work often examines a single way of encoding demographics, overlooking how various signals interact and compound.
+(a) Within-Group Name, $n=5$
+(b) Non-name, $n=5$
+Figure 7: Exclusion metric for retrieval after performing non-demographic perturbations (i.e., within group name changes - left, and modifying spacing and adding typos - right).
 More broadly, do models exhibit brittleness to non-demographic perturbations?
-We consider two sets of perturbation analyses that are non-demographic: (1) How sensitive are models to within-group name perturbations?
+To disentangle fairness from robustness issues, we consider two sets of perturbation analyses that are non-demographic: (1) How sensitive are models to within-group name perturbations?
 and (2) How sensitive are models to non-name perturbations?
-Even when perturbing names within the same demographic group, models surprisingly exhibit similar levels of sensitivity to those observed with demographic name perturbations (Figure 6(a)).
-(a) Within Group Name Perturbation
-(b) Non-name Perturbation
-Figure 6: Exclusion metric for retrieval after performing non-demographic perturbations (i.e., within group name changes - left, and modifying spacing and adding typos - right).
-Results shown for $n=5$.
-For non-name perturbations, we find that models are extremely sensitive to both spacing and typos, although to a lesser extent than name perturbations.
+Even when perturbing names within the same demographic group, models surprisingly exhibit highly similar levels of sensitivity to those observed with gender and race name perturbations (Figure 7(a)).
+We find that models are extremely sensitive to both spacing and typos, but to a lesser extent than names.
 As shown in Figure 7(b), most models demonstrate higher sensitivity to spacing than typos, though there is surprising sensitivity to both.
-In particular, mistral-embed excludes resumes from the top-5 set 72.76% of the time solely based on spacing, which indicates that simple formatting choices can have a massive impact on fairness (in this case, much more than names).
+In particular, mistral-embed excludes resumes from the top-5 set 72.76% of the time solely based on spacing, which indicates that formatting can have a massive impact on fairness (in this case, much more than names).
 Generated resumes display nearly twice the sensitivity to spacing changes compared to web-scraped Kaggle resumes (27.98% vs.
-15.35% averaged across models and $n$ values), likely due to their more structured and formatted nature.
-In summary, we observe that retrieval models lack robustness more generally, which might contribute to fairness-related issues.
+15.35% averaged across models and $n$ values), likely due to their structured formatting.
+In summary, we observe that retrieval models lack overall robustness, which has fairness implications.
 5 Discussion
-Our results highlight important issues about model fairness and brittleness in real-world contexts.
+Our results highlight failures of both fairness and robustness in LLMs in hiring contexts.
+We differ from prior work on LLM fairness in summarization (Zhang et al., 2024) and ranking (Xu et al., 2024) in that our evaluations are grounded in real-world applications, and this reveals novel insights that have ecological validity.
 First, we observe that model rankings in a retrieval setting are impacted considerably by both subtle demographic and non-demographic changes.
-In practice, these differences would lead to unintended exclusion, with candidates being eliminated from consideration at an early stage.
+In practice, these differences would lead to unintended exclusion, with candidates being eliminated from consideration during initial screening.
 We also observe that subtle demographic differences in resumes can alter the way candidates are discussed in generated summaries.
-As a result, even candidates who make it past the initial resume screening stage may be portrayed differently based on demographic attributes.
-These differences can impact downstream decision-making, since recruiters might rely solely on summaries to shortlist candidates.
-Additionally, it is important to consider compositional effects when combining components (i.e., ranking $\rightarrow$ summarization), since biases can compound due to the sequential nature of these tasks.
-Therefore, the candidate pool may (1) leave out qualified candidates through retrieval bias, as well as (2) differentially represent candidates through summarization bias.
-Applying this to our results, at worst case, we find that modifying racial indicators on resumes from W $\rightarrow$ B (using both names and extracurricular information) can result in roughly (1) 70% of candidates being filtered out at the resume screening stage, and (2) 15% of remaining candidates being depicted in less preferable ways.
-Our analysis also reveals that all models are sensitive to non-demographic perturbations.
-These findings suggest that model unfairness may partially stem from more general robustness issues, rather than solely from encoded biases.
-While these insights do not change the real-world impacts of discriminatory outcomes, understanding that disparate treatment can arise from multiple factors—including but not limited to undesirable biases—provides a more complete and nuanced picture for addressing fairness issues.
-Moreover, given that retrieval models are commonly used in RAG systems, these issues may extend to various settings and applications beyond HR.
+As a result, candidates who make it past the initial resume screening stage may be portrayed differently based on demographic attributes, which can impact downstream decision-making.
+Additionally, it is important to consider compositional effects when combining components, since biases can compound due to the sequential nature of these tasks.
+The resulting candidate pool may (1) leave out qualified candidates through retrieval bias, and (2) differentially represent candidates through summarization bias.
+Applying this to our results, in the worst case, we find that modifying racial indicators on resumes (using names + extracurricular information) can result in roughly (1) 70% of candidates being filtered out at the resume screening stage, and (2) 20% of remaining candidates being depicted in less preferable ways.
+Our analysis also reveals that all models are sensitive to non-demographic perturbations, suggesting that model unfairness may partially stem from more general robustness issues, rather than encoded biases alone.
+These perturbations still result in disparate outcomes, but with different underlying causes.
+While these insights do not change discriminatory impact, understanding that disparate treatment can arise from small input changes, both demographic and otherwise, provides a more complete picture for addressing fairness issues.
 6 Conclusion
-We examine fairness in LLM-based hiring systems by analyzing two key components: applicant ranking and summary generation.
-We propose two new metrics to measure fairness in retrieval and generative settings, and find that a hiring pipeline with these components produces biased outcomes, particularly during the initial retrieval phase.
-In addition, we see unexpected sensitivity to minor non-demographic changes, indicating a lack of overall robustness that may contribute to unfair outcomes.
-We hope that our work underscores the need for more real-world evaluations of LLM harms, especially in other high-stakes domains.
+We examine allocational fairness in LLM-based hiring systems by analyzing two key components: applicant ranking and summary generation.
+To support systematic measurement and mitigation of fairness issues, we release a benchmark dataset and introduce an evaluation framework with new metrics.
+[^12] The data and code can be found at: https://github.com/preethisesh/hiring_fairness.
+We find that a hiring pipeline consisting of these two stages produces biased outcomes, particularly during the retrieval phase.
+In addition, models show unexpected sensitivity to minor non-demographic changes, revealing a lack of overall robustness that may contribute to unfair outcomes.
+These findings underscore the need for targeted strategies to improve the fairness of LLM-based hiring, and the importance of realistic, application-grounded evaluations of LLM harms.
 Limitations
-Our analysis focuses exclusively on English resumes and job posts, which poses limitations.
+Our analysis focuses exclusively on English resumes and job posts.
 Future research should investigate fairness considerations in multilingual settings and examine whether our conclusions hold across various languages.
@@
 Understanding these nuances is crucial for evaluating and developing hiring systems that serve diverse global talent pools.
+Since we are releasing our code and datasets, researchers in other regions will be able to expand our work as well.
 While our analysis examines whether hiring systems behave differently for various gender (male and female) and racial (White and Black) groups, it is meant to be illustrative rather than exhaustive and only covers a subset of gender and racial identities.
-We only consider binary gender biases, and exclude non-binary gender biases from our analysis (e.g., if a candidate explicitly declares pronouns on their resume).
-We also only focus on Black and White racial groups, since this is a common emphasis in fairness studies.
-In addition, future research should expand beyond these commonly investigated biases and analyze to what extent other types of demographic information (e.g., age and nationality) impact LLM fairness in hiring.
+We only consider binary gender biases, and exclude non-binary gender biases from our analysis, since this information cannot be inferred from a name.
+While candidates may explicitly declare pronouns on resumes, we do not observe this in the resumes we collect, so we do not vary them.
+In addition, we only focus on Black and White racial groups, since this is a common emphasis in fairness studies, and only to do so in the context of US names.
+We hope future work expands beyond these commonly investigated biases and analyzes the extent to which other types of demographic information (e.g., age and nationality) impact LLM fairness in hiring.
+Moreover, although the way we handle name perturbations is standard practice in NLP fairness literature, we acknowledge that names can encode demographic axes beyond gender and race, including age, class, and region.
+These signals are more subtle and challenging to isolate, making it difficult in practice to vary only a single dimension at a time.
+It is worth noting that we control for other factors such as name frequency to reduce potential confounds.
 Appendix A Appendix
-A.1 Names
+A.1 Additional Background and Related Work
+For background on the allocational fairness of LLMs in high-stakes domains, please see the Introduction.
+Name Perturbations
+Performing name perturbations to study fairness is common practice in NLP fairness literature (Webster et al., 2021; An and Rudinger, 2023; Steen and Markert, 2024; Wan et al., 2023; An et al., 2024).
+We go beyond this by perturbing resumes with extracurricular information, as done in Glazko et al.
+(2024), but largely focus on names because it is common practice.
+It is worth pointing out that Gautam et al.
+(2024) highlight limitations around inferring sociodemographic groups from names, such as poor validity.
+We try to account for some of these concerns by using the carefully curated names from Yin et al.
+(2024).
+Fairness Definitions
+We draw connections between the metrics we use and traditional ML fairness metrics (Mehrabi et al., 2021).
+Non-uniformity is connected to statistical parity, which is satisfied if the probability of a prediction is independent of demographic group.
+We adapt this idea by evaluating for non-uniformity in the demographic distribution of top-x%.
+Exclusion bears resemblance to both individual fairness (Dwork et al., 2012), which assesses whether similar individuals are treated similarly, and counterfactual fairness (Kusner et al., 2017), which assesses whether outcomes are consistent for counterfactual individuals.
+Similarly, exclusion measures the stability of rankings under demographic perturbations.
+Fairness in Summarization and Ranking
+Several studies have identified biases in LLM-generated summaries (Shandilya et al., 2018; Guo et al., 2023; Zhang et al., 2024; Li et al., 2025), but they do not conduct application-grounded evaluations or consider allocational harms.
+A few recent works have also studied the fairness of LLMs in ranking (Wang et al., 2024; Xu et al., 2024).
+Similarly, these works mainly focus on traditional retrieval tasks such as article relevance, rather than real-world LLM usage in high-stakes domains like hiring.
+A.2 Focus on Evaluation over Mitigation
+Given the lack of work on investigating allocational harms in LLM-based hiring systems, our main goal is to establish a comprehensive benchmark of fairness risks.
+Benchmarking is necessary to first understand fairness issues, and mitigation is the natural next step.
+Meaningful progress towards mitigation cannot be made without proper evaluation and metrics—a shared framework is necessary to compare the performance of mitigation methods and track improvements.
+We will make our data and code available to the community, which will enable testing various mitigation approaches.
+A.3 Names
 We use White male, Black male, White female, and Black female names curated by Yin et al. (2024), which we list below:
@@
 Alfreda, Amari, Aniya, Aniyah, Aretha, Ashanti, Ayana, Ayanna, Chiquita, Dasia, Deasia, Deja, Demetria, Demetrice, Denisha, Domonique, Eboni, Ebony, Essence, Iesha, Imani, Jaleesa, Jalisa, Janiya, Kenisha, Kenya, Kenyatta, Kenyetta, Keosha, Keyona, Khadijah, Lakeisha, Lakesha, Lakeshia, Lakisha, Laquisha, Laquita, Lashanda, Lashawn, Lashonda, Latanya, Latasha, Latesha, Latisha, Latonia, Latonya, Latoria, Latosha, Latoya, Latrice, Mahogany, Marquita, Nakia, Nikia, Niya, Nyasia, Octavia, Precious, Quiana, Rashida, Sade, Shakira, Shalonda, Shameka, Shamika, Shaneka, Shanequa, Shanice, Shanika, Shaniqua, Shanita, Shaniya, Shante, Shaquana, Sharita, Sharonda, Shavon, Shawanda, Sherika, Sherita, Tameka, Tamia, Tamika, Tanesha, Tanika, Tanisha, Tarsha, Tawanda, Tawanna, Tenisha, Thomasina, Tierra, Tomeka, Tomika, Towanda, Toya, Tyesha, Unique, Willie, Zaria
-A.2 Professions
+A.4 Resume Dataset Creation and Statistics
+We carefully curate our synthetic resume dataset to systematically vary demographic signals, while still preserving the main content of the resume.
+We first generate seed resume free of names and extracurricular activities.
+Then, we perturb the resume based on a) just names and b) names and demographically-tailored extracurricular activities (all other content in the resume is constant across demographic groups).
+Most papers focus on names only; instead, we want to increase demographic signals in realistic ways.
+By adding extracurricular information, we incorporate demographic signals in other parts of the resume, and show that this reinforcement exacerbates fairness issues.
+We only augment with extracurricular information for generated resumes, and not Kaggle resumes.
+Initially there are 525 generated resumes and 1175 Kaggle resumes [^13] Has a CC0: Public Domain License, without any demographic information.
+For each perturbation type, we then modify the original dataset.
+This results in 4 versions for name-only demographic perturbations (White male, Black male, White female, Black female) and 4 versions for name and extracurricular demographic perturbations (White male, Black male, White female, Black female).
+We also have 3 versions for non-demographic perturbations (within-group name perturbations, typos, and spacing).
+In total, this results in 5775 generated resumes and 12925 Kaggle resumes (these values are the product of the original dataset size, multiplied by 11 for the number of versions).
+A.5 Professions
 We list the professions/fields used in our analysis:
@@
 Apparel, Aviation, Banking, Chef, Construction, Consultant, Finance, Fitness, Healthcare, IT, Teacher
-A.3 Prompts
+In total, there are 22 professions for generated resumes (of which we select 11 for job posts), and 24 fields for Kaggle resumes (of which we select 11 for job posts).
+A.6 Prompts
 Generating Seed Resumes
@@
   • Remove personal information from the resume, and do not generate any names, emails, websites, or addresses.
-  • Do not include any placeholders for personal information (e.g., <PHONE_NUMBER>, <PERSON>, [Name], [Address], [Email], etc.).
+  • Do not include any placeholders for personal information (e.g., [PHONE NUMBER], [Name], [Address], [Email], etc.).
   • The top of the resume should display [PROF].
@@
 Adding Extracurriculars
-You are a/an [RACE], [GENDER] professional.
-Can you add three sections to the resume below: (1) AWARDS, (2) MENTORSHIP AND VOLUNTEERING, and (3) CLUBS AND ORGANIZATIONS, that are reflective of your identity and background?
-If any of these sections (or equivalent sections) are already present, please modify them to better reflect the experiences of a/an [RACE], [GENDER] professional.
-Please output the full revised resume, which includes these additional sections.
-The other sections should remain the same, and please keep the overall format the same as well.
-Keep [NAME] at the top of the resume, and also add ‘——–’ at the top and bottom of the resume.
-Do not output anything else before or after the resume.
+You are [RACE], [GENDER] professional. Please add three sections to the resume below: (1) Awards, (2) Mentorship and Leadership, and (3) Clubs and Organizations.
+  • These additional sections should be reflective of your identity and background.
+  • For each of these sections, output them so that they can be directly added to the resume (i.e., formatted with section headers and bullet points).
+  • Use the exact same section header format (i.e., punctuation, capitalization) present in the resume.
+  • Do not output any introductory or explanatory text. Only output these additional sections.
 Generating Summaries
 Generate a [LENGTH]-word summary of the resume above and its suitability for the role: [PROF]. Make sure the format is a paragraph and written in [POV] person.
-A.4 Example of Extracurricular Information
+A.7 Example of Extracurricular Information
 We present generated extracurricular activities for Latoya Williams vs.
@@
   • Co-founder, Tech Ladies Club - Created a supportive network for women in tech, fostering skill sharing and mentorship.
-A.5 Proxy Measures
+A.8 Proxy Measures
 We use the following measures as proxies for undesirable variation that could influence the decision of an HR staff reading the summary:
   • Reading ease is measured using Flesch Reading Ease score (Kincaid, 1975), with higher scores indicating greater ease.
-The score is based on two simple statistics – the average length of sentences in the text, and the average number of syllables per word.
+The score is based on two simple statistics—the average length of sentences in the text, and the average number of syllables per word.
 [^14] https://pypi.org/project/textstat/
@@
 Figure 8: Human Annotation Results for 3 characteristics (quantification, focus, and individual impact).
-A.6 Human Preferences
+A.9 Human Preferences
 It is unclear whether the chosen measures for summarization (reading ease, reading time, polarity, subjectivity, and regard) capture meaningful differences in summaries.
@@
 We then asked experts [^16] We recruited 6 HR professionals to be annotators (US, Canada, and UK based), and conveyed that annotations would be used towards research on evaluating LLMs in hiring pipelines.
-We did not provide any monetary compensation to annotate the preferred summary in the pair (200 pairs annotated in total), and investigated whether experts displayed consistent preferences with respect to the characteristics being varied (quantification, focus, and individual impact).
+We did not provide any monetary compensation.
+to annotate the preferred summary in the pair (200 pairs annotated in total), and investigated whether experts displayed consistent preferences with respect to the characteristics being varied (quantification, focus, and individual impact).
 We gave the following instructions:
@@
 Which resume summary below do you prefer?
-We find that 4 out of 6 annotators favor the use of quantification, while 1 annotator prefer no quantification (Appendix Figure 7(a)).
+We find that 4 out of 6 annotators favor the use of quantification, while 1 annotator prefers no quantification (Appendix Figure 8(a)).
 We see that 4 out of 6 annotators demonstrate a modest preference for focus, with the other 2 remaining neutral (Appendix Figure 8(b)).
@@
 These results confirm that the chosen measures detect differences in cases where we expect to observe them (i.e., based on results from human preferences).
+A.10 Summarization Fairness Metric
+To measure fairness in summarization, we compute invariance violations, which computes the percentage of t-tests for which the null hypothesis is rejected.
+The total number of t-tests corresponds to $M\times A\times C\times T\times L\times P$, where
+  • $M$: # $\text{\# of models}=6$
+  • AA: # $\text{\# of automated measures}=5$
+  • $C$: # $\text{\# of demographic comparisons}=4$
+  • $T$: # $\text{\# of temperature settings}=2$
+  • $L$: # $\text{\# of length settings}=2$
+  • $P$: # $\text{\# of point-of-view (POV) settings}=2$
+When computing invariance violations, we group or aggregate results to get a percentage for each model and demographic comparison type (gender, which considers MW-FW and MB-FB comparisons, and race, which considers MW-MB and FW-FB comparisons).
+Within each group, we perform Benjamini-Hochberg correction (Benjamini and Hochberg, 1995) to account for multiple comparisons.
+These results are shown in Figure 2.
+We also perform Bonferroni correction (Bland and Altman, 1995) as an alternate method to address multiple comparisons, which is more aggressive in its correction of false positives.
+We show results using this method in Figure 9.
+(a) Gender
+(b) Race
+Figure 9: Summarization Results: Invariance violations for generated summaries, separated by completion model and perturbation type.
+Results are shown across 5 runs.
+Left 3 models are considered "smaller" models, right 3 models are considered "larger" models.
+Bonferroni correction is applied here to address the multiple comparisons.
+A.11 Retrieval Fairness Metrics
+We compute two retrieval fairness metrics: non-uniformity, introduced by Wilson and Caliskan (2024) and exclusion, which we propose.
+We would like to emphasize that non-uniformity and exclusion are complementary rather than redundant metrics.
+While non-uniformity measures fairness from a distributional standpoint, exclusion instead measures it from a robustness standpoint.
+Intuitively, they answer different questions about fairness in retrieval:
+Non-uniformity
+Let us consider demographically perturbed but otherwise equivalent resumes for four demographic groups: Black female, White female, Black male, and White male.
+Non-uniformity answers the question: are the four groups represented unequally in the top-xx% of retrieved resumes.
+Exclusion
+Let us consider the top-$n$ White male resumes for a given job post.
+Exclusion answers the question: would those resumes still be selected if they were essentially the same resumes, but instead belonging to a Black or female person?
+As we see in Figures 3 and 4, the two metrics lead to different conclusions about the best retrieval model in terms of fairness (text-embedding-3-small for non-uniformity vs.
+embed-english-v3.0 for exclusion).
+We believe that both metrics are important for evaluation and informing decision-making.
+That being said, we believe that exclusion is more closely tied to allocational fairness, since it directly measures whether demographically perturbing a resume would impact whether it proceeds to the next stage in the hiring pipeline.
+A.12 Justifying Chosen Non-demographic Perturbations
+We consider two non-demographic perturbations: spacing and typos.
+We expect both formatting and typos to have minimal impact on an embedding-based retrieval system, since embedding models are trained on noisy web text and do not have explicit resume supervision data.
+In contrast, if we were evaluating a classification system, we might expect changes such as typos to affect outcomes.
+Our goal in applying non-demographic perturbations is to establish a meaningful comparison point for studying the impact of demographic perturbations.
+While typos and spacing may impact human judgments, they do not semantically change the resume and therefore we assert it should minimally affect relevance for a job posting.
+Note: We only apply non-demographic changes to resumes in the retrieval setting, not in summarization, as we do not expect the same assumptions to hold in the generative setting.
 (a) Top-5%
@@
 Figure 10: Non-uniformity metric for top-5 and top-10% of retrieved Kaggle resumes.
-Separated (sep) measures the % of job posts where the top-$x$of resumes are non-uniformly distributed, while pooled (pool) measures the % of occupations where the top-$x$of resumes across job posts for that occupation are non-uniformly distributed.
-A.7 Retrieval Fairness Metrics
-We would like to emphasize that non-uniformity and exclusion are complementary rather than redundant metrics.
-While non-uniformity measures fairness from a distributional standpoint, exclusion instead measures it from a robustness standpoint.
-Intuitively, they answer different questions:
-Non-uniformity
-Let us consider demographically perturbed but otherwise equivalent resumes for four demographic groups: Black female, White female, Black male, and White male.
-Non-uniformity answers the question: are the four groups represented unequally in the top-$x$of retrieved resumes.
-Exclusion
-Let us consider the top-$n$ White male resumes for a given job post.
-Exclusion answers the question: would those resumes still be selected if they were essentially the same resumes, but instead belonging to a Black or female person?
-As we see in Figures 2 and 3, the two metrics lead to different conclusions about the best retrieval model in terms of fairness (text-embedding-3-small for non-uniformity vs.
-embed-english-v3.0 for exclusion).
-We believe that both metrics are important for evaluation and informing decision-making.
+Separated (sep) measures the % of job posts where the top-xx% of resumes are non-uniformly distributed, while pooled (pool) measures the % of occupations where the top-xx% of resumes across job posts for that occupation are non-uniformly distributed.
 (a) Gender, $n=5$
@@
 M→F perturbs male names to female names and F→M perturbs female names to male names, while W→B perturbs White names to Black names and B→W perturbs Black names to White names.
-(a) Within-group Name, $n=5$
-(b) Within-group Name, $n=10$
-(c) Within-group Name, $n=100$
+(a) Within-group, $n=5$
+(b) Within-group, $n=10$
+(c) Within-group, $n=100$
 (d) Non-name, $n=5$
@@
 M→F perturbs male names to female names and F→M perturbs female names to male names, while W→B perturbs White names to Black names and B→W perturbs Black names to White names.
-(a) Within-group Name, $n=5$
-(b) Within-group Name, $n=10$
-(c) Within-group Name, $n=100$
+(a) Within-group, $n=5$
+(b) Within-group, $n=10$
+(c) Within-group, $n=100$
 (d) Non-name, $n=5$